### v1

# An Entropy-based Text Watermarking Detection Method

Abstract
Currently, text watermarking algorithms for large language models (LLMs) can embed hidden features to texts generated by LLMs to facilitate subsequent detection, thus alleviating the problem of misuse of LLMs. Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved. In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token should be adjusted according to its entropy during watermark detection, rather than setting the weight of all tokens to the same value as in previous methods. Specifically, we proposed an Entropy-based Watermark Detection (EWD) that gives higher-entropy tokens higher weights during watermark detection, so as to better reflect the degree of watermarking. Furthermore, the proposed detection process is training-free and fully automated. In the experiment, we found that our method can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions. Our code and data will be available online.
An Entropy-based Text Watermarking Detection Method
Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1 1The Chinese University of Hong Kong 2Tsinghua University
1 Introduction
Figure 1: Watermarking algorithm has limited effect on a low-entropy token as compared to a high-entropy token. Therefore, a low-entropy token has a much smaller probability of being watermarked. In a code generation scenario, most of the tokens are low-entropy, resulting in less watermarked tokens and a small detection z-score.
The advancements in LLMs have enabled it to generate high-quality outputs indistinguishable from humans, and also achieve better performance on various real-world generation tasks. However, this gives rise to the potential risk of misuse. Specifically, LLM-generated assignments, including essays, code, and reports, pose a threat to academic integrity. Misusing LLMs to generate fake news can also lead to negative societal consequences Megías et al. (2021). Therefore, algorithms capable of effectively marking and identifying machine-generated contents have become essential.
The text watermarking algorithm for LLMs can embed hidden features that can be detected later during the text generation process of LLMs, thus effectively alleviating the problem of misuse of LLMs. The current LLM watermarking technology is usually to embed watermark information in the process of generating tokens during LLM inference. For example, Kirchenbauer et al. (2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list. By modifying the tokens, it leads the generated text to contain mostly watermarked (green) tokens. In the watermark detection process, the degree of watermark is usually judged by detecting the number of watermarked tokens in the text.
Although the current text watermarking algorithm performs well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved. The reason behind this is that low-entropy tokens usually have stronger determinism, are more difficult to modify in the watermarking process, and are easier to affect the quality of the text when modified.
In figure 1, we illustrate the watermark ratio in tokens of different entropy levels. A token with higher entropy will have a significantly larger probability to be modified and watermarked. On the contrary, a low-entropy token has a much lower probability.
When the generated text contains a large number of low-entropy tokens, tokens that can hardly be modified by the watermarking algorithm, the detection results of the watermark detector are more easily affected by the state of these unmodified low-entropy tokens. For example, the low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments. To address this problem, Lee et al. (2023) proposed an entropy-based selective watermark detection method named SWEET. Specifically, they only detect tokens that exceed a certain entropy threshold, thus excluding tokens that are difficult to be modified by the watermark generator. However, this method also has some problems. First, their method requires a manually set entropy threshold, which needs to be determined by analyzing samples of code written by humans. At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
To address the aforementioned problems on low-entropy watermark detection, we introduced an Entropy-based Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text. Specifically, we argue that the importance of a token during detection should be proportional to its entropy. For example, the low-entropy tokens should have a limited impact on the detection result. Therefore, we assign an importance weight to each generated token based on its entropy, which is continuous rather than binary as in SWEET. Then, inside the detector, the watermark level of a given text will be the weighted sum of all the watermarked tokens. By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm failed to watermark a low-entropy token, it will make minimal impact to the detected watermark level. The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations. Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy level. There is no need to switch between different detectors facing various kinds of input texts.
In the experiment, we evaluated the detection accuracy of our proposed methods in various low-entropy generation tasks, including code generation. We also test the versatility of our method in text of high entropy level, such as document-to-text generation. Overall, our method outperformed baselinse methods including KGW and SWEET in detection accuracy.
In summary, the contributions of our work are summarized as follows:
  • We proposed an entropy-based watermark detection method called EWD.
  • We provide a theoretical analysis of the detection accuracy in our method in a low-entropy setting.
  • Through experiments, we show that our method not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
2 Related Work
Watermark message can be embedded into existing machine-generated text by performing a series of modifications, such as lexical substitutions Munyer and Zhong (2023); Topkara et al. (2006); Yang et al. (2022); Yoo et al. (2023a); Yang et al. (2023). Those methods typically replace words with their alternatives to place watermark. For example, Topkara et al. (2006) used WordNetFellbaum (1998) as his dictionary while Munyer and Zhong (2023) utilized a pretrained Word2Vec model to select the alternative words. To ensure that the repalced words would cause minimal changes to the text semantics, context-aware lexical substitution was introduced. For example, Yang et al. (2022) proposed a BERT-based infill model to generate replacement with regard to the sentence meaning. To improve the robustness of the algorithm against possible corruptions, Yoo et al. (2023a) fine-tuned the infill model to generate consistent word distribution given several corrupted texts of the same origin. However, the level of modification to an existing text is often limited due to the need to protect text quality, so the capacity and effectiveness of lexical-based watermarks are largely restricted.
Watermark embedding can also be conducted during the LLM generation phase, either by modifying the token sampling or model logits. Christ et al. (2023) introduced a technique that incorporates watermark message by predefining a random number sequence for sampling the tokens. Detection process is to check the correlation of the input text with random number sequences. To enhance resistance to text modifications, Kuditipudi et al. (2023) utilized the Levenshtein distance in matching the text with number sequences. For the latter, the most representative one that modifies the model logits is proposed by Kirchenbauer et al. (2023a) based on previous tokens during text generation. Specifically, the model vocabulary will be separated into green and red lists during each generation step. And the logit value of every green token will be enhanced by a positive constant. There are a number of works proposed on top of this algorithm with enhanced performance in payload Yoo et al. (2023b); Wang et al. (2023), robustness Zhao et al. (2023); Liu et al. (2023b); Kirchenbauer et al. (2023b); Ren et al. (2023) and unforgeabilityLiu et al. (2023a); Fairoze et al. (2023). For example, Yoo et al. (2023b) extends the method to a multi-bit setting to allow the embedding of longer messages. Wang et al. (2023) proposed to utilize a proxy LLM for multi-bit message embedding. To increase robustness against editing attacks, Zhao et al. (2023) proposed to set a universal green and red lists during the text generation process. Liu et al. (2023b) proposed a different approach which utilizes the semantic meaning of previous generated text to determine the green and red lists. Also, to enhance the unforgeability of text watermark, Liu et al. (2023a) introduced a neural network for text generation and detection. In this work, we focus more on the detection of low-entropy texts. Lee et al. (2023) designed a selective watermark generation and detection approach for low-entropy code generation. However, it requires a human code dataset to help manually determine the entropy threshold.
3 Preliminaries
3.1 LLM
Here we introduce the necessary concepts used in this paper. A large language model (LLM) takes a prompt as input and outputs subsequent tokens as the corresponding response. Specifically, we denote the initial prompt as $x^{prompt}$. Then, at the $l$-th step, the input to the LLM is the $x^{prompt}$ and the already-generated tokens $T_{:(l-1)}$. The LLM would generate a distribution logits $P_{LLM}(x^{prompt},T_{:(l-1)})$, consisting of $P_{LLM}(v|x^{prompt},T_{:(l-1)})$ over each token $v$ of its model vocabulary based on the joint input. The token sampling is based on the logits and specified sampling strategies, such as greedy sampling.
3.2 Text Watermarking
Text watermarking is to embed a piece of message into the given text that can be decoded for the purpose of identification. Here we include a specific branch of it, defined as Watermark for LLMs by Liu et al. (2024), specifically the ones concerning Logits Generation. The general watermarking scheme used in this paper is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$. The modification on the original logits by adding a small value to the logits of green-list tokens will result in LLM’s preference on those tokens in the output text. Thus such modification could likely produce a watermarked (green) token. The detection algorithm is to calculate the watermark level using the following equation
$z=(|s|_{G}-\gamma T)/\sqrt{\gamma(1-\gamma)T},$ (1)
where $|s|_{G}$ is the number of watermarked tokens and $\gamma$ is the green-list ratio in the model vocabulary. The algorithm would output 1 if the value is significantly greater than a given threshold, and 0 otherwise. In this work, we only concern the watermark detection process.
3.3 Token Entropy and Low-entropy Scenario
For the logits during each generation step, its degree of spread-out can be measured by calculating the entropy $SE(P_{LLM}(x^{prompt},T_{:(l-1)}))$. In this paper, we follow the spike entropy from Kirchenbauer et al. (2023a),
$SE(k)=\sum_{k}{\frac{p_{k}}{1+zp_{k}}},$ (2)
where $p_{k}$ is the probability vector of the token $k$ and $z$ is a scalar. The entropy value achieves minimal if the logits concentrates on a single token and maximal given a uniform distribution.The watermark generation process is more effective on a high-entropy token where the logits distribution is nearly uniform, so that the modification could produce a green-list token easily. On the contrary, a low-entropy token is almost deterministic and hard to be modified. Therefore, the watermarked ratio for a low-entropy token is much lower, resulting in fewer watermarked tokens in the output. A low-entropy scenario refers to the LLM-generated output consisting of a great number of low-entropy tokens. Here we denote the low-entropy token ratio to the whole text as $\varepsilon$.
4 Method
In this section, we will first explain the motivation by analysing the watermark effect on low-entropy tokens and previously-proposed detection methods, then provide a detailed introduction to the proposed EWD algorithm.
4.1 Motivation
In KGW Kirchenbauer et al. (2023a)’s watermark generation method, a positive constant is added to the logits value of all green-list tokens so that the final output would prefer such tokens. The distribution level of the logits, represented by the entropy, matters significantly during this process. When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this watermark embedding process would very likely produce a green-list (watermarked) token. An average-high-entropy text would then consist of many watermarked tokens and output a high z-score according to Eq.1. However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would produce a watermarked token with a much lower and constant probability. As a result, the number of watermarked tokens and the consequence z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1.
4.2 Proposed Method
As discussed in the previous sections, it is crucial for the detector to consider each token’s entropy. And we take a step further than SWEET: assigning an entropy-based weight to each token, representing their significance in the detection process.
In figure 1, we can observe that a high-entropy token has a larger probability of being watermarked or is easier to be modified and watermarked than a low-entropy token. Since a low-entropy token is hard to modify, then whether it is watermarked or not should have a limited impact on the overall watermark level. Specifically, we argue that when an easy-to-watermark token is not watermarked, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-modify token. That is to say, The impact of a non-watermarked token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its likelihood to be modified, and same for the impact of a watermarked token to yielding a positive detection result:
$I(Detect(T)=\textbf{1}(t\in G)|t)\propto Entropy(t).$ (3)
When reflected in the real world, this impact is represented by the token weight. Token with a larger spike entropy, meaning that it is easier to be modified, will be given a larger weight. On the contrary, a low-entropy token will be given a much lower weight close to 0. As the sum of all watermarked token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a non-watermarked low-entropy token than a non-watermarked high-entropy token, and a larger gain when observing a watermarked high-entropy token than a watermarked low-entropy token.
Algorithm 1 Entropy-based Watermark Detection
Input: a language model $M$, input text $T$, detection key $k$, window size $m$, detection threshold $\tau$, green list ratio $\gamma$
Compute token entropy based on model logits, $SE=SpikeEntropy[M(t)]=[SE_{0},SE_{1}....]$
Compute token weights based on token entropy, $W=ComputeWeight[SE]=[W_{0},W_{1}....]$
for $l=m,m+1....$ do
     Use the detection key $k$ and previous $m$ tokens $T_{l-m:l-1}$ to find the green list $G$
     If current token $T_{l}$ is inside the green list $G$, then add its weight $W_{l}$ to $|s|_{G}$
Compute the $z$-statistic:
$\centering z=(|s|_{G}-\gamma\sum_{i=m}^{|T|-1}{W}_{i})/\sqrt{\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}}\@add@centering$ (4)
if $z>\tau$ then return 1, $i.e.,$Input text is watermarked"
else return 0, $i.e.,$Input text is not watermarked"
Pseudocode of our proposed detection method is provided in Algorithm 1. We first acquire the text entropy by computing the model logits of each token. The next step is to determine the importance weight of each token, the output of a designed function with their entropy values as inputs. Due to the narrow range of entropy values, our goal here is to create a bigger gap between tokens of high entropy and low entropy. We achieve this by the function $ComputeWeight$, which is training-free and a totally automated process. Here, we first normalize the text entropy. Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights:
$W=Linear(Norm(SE)).$ (5)
After obtaining the weights for each token, we then apply the standard detection procedure by Kirchenbauer et al. (2023a) to determine the group of watermarked tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens. Finally, we sum up the weights of the watermarked tokens and calculate the z-statistic using Eq.4 given the sum and weight vector of all detected tokens. The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
In real-world low-entropy scenarios, such as code generation, there are a large number of hard-to-modify tokens. When detecting watermarked texts with our proposed EWD, the vast majority of non-watermarked low-entropy tokens will not prevent the detector from yielding a positive conclusion. Meanwhile, facing non-watermarked texts, any non-watermarked high-entropy token will strongly suggest the detector to make a negative output.
5 Theoretical Analysis
In this section, we theoretically analyzed the Type-I and Type-II error of our proposed detection method. Specifically, we compared EWD with two previous methods, KGWKirchenbauer et al. (2023a) and SWEETLee et al. (2023).
To do so, we adopted Kirchenbauer et al. (2023a)’s steps of approach. Each token in both watermarked and non-watermarked texts can either be green or not, so we can model it as a binomial distribution. Therefore, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes. Finally, we calculate and compare the probability of the sum surpassing or falling behind the detection threshold $|\tilde{s}|_{G}$, which is the $|s|_{G}$ value in each detection function with a fixed $z$.
5.1 Type-I Error
Type-I error measures the likelihood a human-written text is incorrectly identified as being generated by a watermark algorithm.
In a human-written text, each token is independent of the watermark algorithm. Therefore, its probability to be included in the green list is $\gamma$. Therefore, KGW Kirchenbauer et al. (2023a)’s sum of green tokens is approximated by $\mathcal{N}(\gamma T,\gamma(1-\gamma)T)$. For SWEETLee et al. (2023) that only considers high-entropy tokens, such distribution is modelled as $\mathcal{N}(\gamma\tilde{T},\gamma(1-\gamma)\tilde{T})$, where $\tilde{T}$ is the text length excluding low-entropy tokens. Our method EWD added weights to tokens based on entropy, leading to a weighted binomial distribution. We could also approximate it to a normal distribution with mean $\mu=\gamma\sum_{i=m}^{|T|-1}{W}_{i}$ and variance $\sigma^{2}=\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}$.
When $z$ is set to 2, the probabilities for $|s|_{G}$ under each method to surpass its detection threshold $|\tilde{s}|_{G}$ is identically 2.28%.
Table 1: Main results on different datasets using KGW, SWEET and our method for detection.
Code Detection (Low-entropy Scenario)
HumanEval MBPP
Methods 1%FPR 5%FPR Best 1%FPR 5%FPR Best
TPR F1 TPR F1 F1 TPR F1 TPR F1 F1
KGW 0.383 0.551 0.519 0.663 0.776 0.113 0.201 0.327 0.477 0.744
SWEET 0.662 0.793 0.744 0.835 0.897 0.482 0.647 0.692 0.795 0.876
EWD 0.684 0.809 0.850 0.897 0.900 0.615 0.757 0.748 0.833 0.888
Rotowire (High-entropy Scenario)
Sampling Beam-search
Methods 1%FPR 5%FPR Best 1%FPR 5%FPR Best
TPR F1 TPR F1 F1 TPR F1 TPR F1 F1
KGW 0.986 0.988 0.993 0.972 0.990 1.000 0.996 1.000 0.976 0.999
SWEET 0.986 0.988 0.993 0.972 0.990 0.997 0.994 1.000 0.976 0.996
EWD 0.984 0.987 0.997 0.974 0.988 0.999 0.995 1.000 0.976 0.997
C4 (High-entropy Scenario)
Sampling Beam-search
Methods 1%FPR 5%FPR Best 1%FPR 5%FPR Best
TPR F1 TPR F1 F1 TPR F1 TPR F1 F1
KGW 0.998 0.995 0.999 0.976 0.998 0.998 0.995 0.999 0.976 0.998
SWEET 0.995 0.993 0.997 0.975 0.993 0.995 0.993 0.997 0.975 0.993
EWD 0.996 0.993 0.997 0.975 0.996 0.996 0.993 0.997 0.975 0.996
5.2 Type-II Error
Type-II error measures the likelihood a machine-generated text is incorrectly identified as human-written.
For the purposes of analysis, we assumed that a sequence generated by the watermark algorithm has a length of $T=200$ tokens and that the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
Originally, Kirchenbauer et al. (2023a) sampled over 500 watermark-generated new scripts and reported an average spike entropy of 0.807. In this section, we focused specifically on cases with a much lower average spike entropy. Based on our data, we approximated the distribution of spike entropy with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$. Under this distribution, the average spike entropy is 0.607. The following probability from Kirchenbauer et al. (2023a) showed a lower bound of the probability of a token $k$ being watermarked by a watermark generator:
$\mathbb{P}[k\in G]\geq\frac{\gamma\alpha}{1+(\alpha-1)\gamma}SE(k),$ (6)
where $\alpha=exp(\delta)$. We would utilize this probability later in the calculation of $\mu$ and $\sigma^{2}$.
KGW. Based on Eq.6, the statistical $\mu$ and upper bound of $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]$ and $\sum_{k}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. When the average entropy $SE^{*}$ given by the simulated power law distribution is 0.607, we approximated $|s|_{G}$ by $\mathcal{N}(106.9,7.1)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 114) for detection, the false negative rate exceeds 84.6%, or in other words, the detection rate of watermarked text is less than 15.4%. This means that KGW is not good at detecting the watermarked texts in cases of low entropy.
SWEET. Following the detection scheme in SWEET, we set the entropy threshold at 0.695 and excluded all tokens below this threshold during detection. The average spike entropy $SE^{*}$ and remaining text length $\tilde{T}$ then changed to 0.82 and 24 respectively. Using the equations from KGW, we can get the approximation as $\mathcal{N}(17.5,2.2)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 17.2) for detection, the false negative rate exceeds 42%, or in other words, the detection rate of watermarked text is around 58%.
EWD. We assigned higher weights to tokens with high spike entropy in order to achieve a better detection rate in cases of low average spike entropy. Note that in experiments, we use several weight functions, and here, for analysis purposes, we used a linear weight function $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant. Based on Eq.6, the statistical $\mu$ and $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ and $\sum_{k}{W(k)}^{2}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$. Therefore, we can approximate the $|s|_{G}$ detected by EWD with $\mathcal{N}(5.79,0.56)$. If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.3%, or in other words, the detection rate of watermarked text is around 66.7% (Proof and details in Appendix A). Overall, Our method has a type-II error of 33.3%, much lower than KGW and SWEET. Empirical experiments also demonstrate similar results.
6 Experiments
6.1 Experiment Settings
Figure 2: The left figure shows the z-scores of watermarked and human texts in the Rotowire, HumanEval and MBPP datasets respectively, each being detected with 3 different methods. The right figure shows the relationship between token weights and the probability of being watermarked in both watermarked and human texts.
Figure 3: We utilized a variety of weight functions and measured their performance in the code detection datasets. The left figure is a visualization of all studied functions, with normalized spike entropy as input. The right figure shows each function’s detection F1 under 5% FPR with comparison to the SWEET baseline. Each data point can correspond to the illustrated function on the left figure by looking at its shape and x-axis value.
Tasks and Datasets. First of all, we used the KGW Kirchenbauer et al. (2023a) for watermark generation in all the following tasks.
To evaluate the detection performance of our proposed method, we targeted at a commonly-seen low-entropy scenario, code generation. We adopted two datasets following Lee et al. (2023), HumanEval Chen et al. (2021) and MBPP Austin et al. (2021). We used StarCoder Li et al. (2023) for code generation. And we used the human-written answers as non-watermarked texts. During evaluation, we restricted the minimal length of both samples to be 15 tokens.
Besides code generation, we fine-tuned a T5-small Raffel et al. (2023) model for data-to-text generation. Specifically, we utilized Wu et al. (2022)’s version of the Rotowire dataset Wiseman et al. (2017), which was converted for a text-to-table task. The dataset consists of NBA game records, presented as tables, and human-written descriptions. We used the table information as input and asked the fine-tuned T5 to generate descriptive texts by binomial sampling and beam-search. For each input prompt, the models are expected to generate 200 ± 5 tokens.
We also utilized the pre-trained OPT-1.3B Zhang et al. (2022) for news report generation. Our input prompts are from C4 Raffel et al. (2020) dataset. We restricted the maximum prompt length to 50 tokens. For each input prompt, the models are expected to generate 200 ± 5 tokens using binomial sampling and beam-search.
Baselines and Evaluation Metrics. We selected two detection methods as the baselines to compare with. The first is KGW Kirchenbauer et al. (2023a), which is the original detector. The second is SWEET Lee et al. (2023), which excludes low-entropy tokens during detection. We follow the instructions in SWEET to select entropy threshold, namely from the median, mean, and 3rd quartile (Q3) values of the entropy computed from human texts. For evaluation, we set the false positive rate (FPR) under 5%, following Zhao et al. (2023)’s instruction. We reported the true positive rate (TPR) and F1 score for each method.
Hyper-parameters. For watermarked generation, we set $\gamma$ and $\delta$ to 0.5 and 2 respectively. In text generation using binomial sampling, we set the sampling temperature to 0.7. In text generation using beam-search, we set the number of beams to 8. In order to avoid excessive repetition of text, we set "no_repeated_ngrams" to 16.
6.2 Main Results
Before detection, we measured the generation quality of all generated texts. Specifically, we are able to achieve 31.0% and 30.4% in pass@1Chen et al. (2021) for HumanEval and MBPP dataset respectively, comparable to Lee et al. (2023)’s. For the Rotowire dataset, our generated output can achieve 7.24 in text perplexity (PPL). For the C4 dataset, our generated output can achieve 3.25 in PPL. PPL is calculated using OPT-6.7B Zhang et al. (2022).
Table 1 shows that for the low-entropy scenario, code detection, our detector would achieve the best detection accuracy against other baselines. Specifically, for HumanEval and MBPP respectively, our method could achieve a 10.6% and 5.6% improvement than SWEET in true positive rate while maintaining a low-than 5% false positive rate.
Table 1 shows that for the high-entropy scenarios using Rotowire and C4 datasets, our method is better than SWEET and is very similar to the original KGW method, with very minor differences.
6.3 Analysis
Impact of Entropy-based Weights. In figure 2, we give a few visualizations to show the reason why our method surpassed other baselines in detection performance. Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods. Our detection methods would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts. The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset. By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts. In figure 2(b), we plotted the relationship of entropy-based weight and probability of being watermarked. In watermarked text, the probability is proportional to the token weight. That is to say, the tokens assigned with a large weight are very likely to be watermarked and contribute to a larger z-score. On the contrary, in human-written texts, there doesn’t exist such a relationship, explaining why the z-scores of human text remain low.
Ablation Study. In the experiments, we utilized a linear function to gain the entropy-based weights for each token. We also studied a variety of monotonic increasing functions as shown in figure3. Specifically, each function’s gradient is different. For our custom Sigmoid functions, the gradient starts off high and gradually decreases, that is to say, tokens with relatively low entropy will have significant differences in weights, while tokens with relatively high entropy will exhibit smaller differences. Similarly for our custom Exponential functions, the gradient starts off low and gradually increases, meaning that tokens with relatively low entropy will have smaller differences in weight, while tokens with relatively high entropy will exhibit larger differences. We use function strength to represent how far the function deviates from being linear, for example, "Exp 10" deviates from linear more than "Exp 8".
In the experiments with those functions and two code generation datasets, we can observe that the total 8 listed functions could surpass the SWEET baseline in terms of detection F1. This reflects the versatility of our method, as it is applicable to a wide range of functions.
Time Complexity.
Table 2: The table below shows the average time taken to generate 200-token texts using KGW generator, as well as the average time taken for detection using different methods, measured in seconds (s). The experiment is conducted on a NVIDIA RTX A6000 GPU.
Method Generation Detection
KGW 0.0391
SWEET 9.489 0.0830
Ours 0.0831
Compared to the KGW detection, both SWEET and our proposed EWD would demand extra time to compute token entropy. Specifically, our detector would also compute weight given the token entropy before calculating the final z-score. According to table 2, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our detector respectively. In spite of being double the value of KGW, our detector remains highly efficient, and the difference with SWEET is virtually insignificant.
7 Conclusion
In this work, we proposed a text watermark detection algorithm called EWD based on token entropy. We gave theoretical analysis on the detection properties of our method in low-entropy scenarios with comparison of previous works. We also validated the detection performance of our method in the low-entropy code generation task with a number of weight functions. Furthermore, our method would achieve similar performance in high-entropy scenarios as the previous works. Experiments also showed that compared to previous methods, our method would not bring extra computational cost.
Limitations
The limitations of our method mainly include two parts. The first part is that the low-entropy datasets tested are limited. We only conducted experiments on two code generation datasets, and we did not conduct experiments on more low-entropy datasets. At the same time, our method should be theoretically effective for various types of watermark methods, such as token sampling-based methods. This paper did not try more types of methods, which is an important future work direction.
Appendix A Proof of Type-II Error for EWD
Here we demonstrate how we obtained the theoretical Type-II error for EWD. First of all, we attempted to calculate the theoretical $\mu$ and $\sigma^{2}$ based on Eq.6. The theoretical $\mu$ is the weighted sum of green tokens in the watermarked text, using ${|s|_{G}}_{k}$ to represent the weighted number of green tokens from each token $k$,thus we can obtain:
$\mu$ $=\sum_{k}{|s|_{G}}_{k}$ (7)
$\mu$ $=\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ (8)
$=T\mathbb{E}\frac{\gamma\alpha SE}{1+(\alpha-1)\gamma}(SE-C_{0})$ (9)
$=T\frac{\gamma\alpha}{1+(\alpha-1)\gamma}(\mathbb{E}[SE^{2}]-C_{0}\mathbb{E}[SE]).$ (10)
Given the simulated power law of spike entropy, we can obtain $E[SE^{2}]$ and $E[SE]$ as 0.377 and 0.608 respectively, producing $\mu$ equal to 5.79. And for $\sigma^{2}$, we first obtain the variance for each token $k$:
${\sigma_{k}}^{2}$ $=\mathbb{E}[({|s|_{G}}_{k}-\mathbb{E}[{|s|_{G}}_{k}])^{2}]$ (11)
$=\mathbb{E}[{|s|_{G}}_{k}^{2}]-\mathbb{E}[{|s|_{G}}_{k}]^{2}$ (12)
$=\mathbb{P}[k\in G]\cdot W(k)^{2}-\mathbb{P}[k\in G]^{2}\cdot W(k)^{2}$ (13)
$=W(k)^{2}\cdot(\mathbb{P}[k\in G])(1-\mathbb{P}[k\in G]).$ (14)
Therefore, if we represent $\frac{\gamma\alpha}{1+(\alpha-1)\gamma}$ by $C_{1}$, the variance of the whole text can be formulated as:
$\sigma^{2}$ $=\sum_{k}W(k)^{2}\cdot(\mathbb{P}[k\in G])(1-\mathbb{P}[k\in G])$ (15)
$=T\mathbb{E}[W(k)^{2}C_{1}SE(1-C_{1}SE)]$ (16)
$=TC_{1}\mathbb{E}[(SE-C_{0})^{2}SE(1-C_{1}SE)]$ (17)
$=TC_{1}(-C_{1}\mathbb{E}[SE^{4}]+(2C_{0}C_{1}+1)\mathbb{E}[SE^{3}]$
$\ -(2C_{0}+{C_{0}}^{2}C_{1})\mathbb{E}[SE^{2}]+{C_{0}}^{2}\mathbb{E}[SE]).$ (18)
Given the simulated power law of spike entropy, we can obtain $E[SE^{4}]$ and $E[SE^{3}]$ as 0.24 and 0.158 respectively, producing $\sigma$ equal to 0.56. When the detection threshold $z$ is set as 2 (corresponding to $|\tilde{s}|_{G}$ = 5.55), the probability of $|s|_{G}\sim\mathcal{N}(5.79,0.56)$ falling below 5.55 is 33.3%, $i.e.$ Type-II error.

### v2

@@
 Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved.
-In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token should be adjusted according to its entropy during watermark detection, rather than setting the weight of all tokens to the same value as in previous methods.
+In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token during watermark detection should be adjusted according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods.
 Specifically, we proposed an Entropy-based Watermark Detection (EWD) that gives higher-entropy tokens higher weights during watermark detection, so as to better reflect the degree of watermarking.
@@
 (2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list.
-By modifying the tokens, it leads the generated text to contain mostly watermarked (green) tokens.
-In the watermark detection process, the degree of watermark is usually judged by detecting the number of watermarked tokens in the text.
-Although the current text watermarking algorithm performs well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved.
-The reason behind this is that low-entropy tokens usually have stronger determinism, are more difficult to modify in the watermarking process, and are easier to affect the quality of the text when modified.
-In figure 1, we illustrate the watermark ratio in tokens of different entropy levels.
+By modifying the token distributions, it leads the generated text to contain mostly watermarked (green) tokens.
+In the watermark detection process, the degree of watermark is usually judged by the number of watermarked tokens in the text.
+Although current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved.
+The reason behind this is that low-entropy tokens usually have stronger determinism, are more difficult to modify in the watermarking process, and are easier to affect the quality of the text when forcefully modified Tu et al.
+(2023).
+In figure 1, we illustrate the watermark ratio in tokens of three different entropy levels.
 A token with higher entropy will have a significantly larger probability to be modified and watermarked.
-On the contrary, a low-entropy token has a much lower probability.
-When the generated text contains a large number of low-entropy tokens, tokens that can hardly be modified by the watermarking algorithm, the detection results of the watermark detector are more easily affected by the state of these unmodified low-entropy tokens.
-For example, the low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments.
+On the contrary, it is rare to see a green token in low-entropy tokens.
+When the generated text contains a large number of low-entropy tokens, tokens that can hardly be modified by the watermarking algorithm, the detection results of the watermark detector are more easily affected by the states of these unmodified low-entropy tokens even when the watermark generator managed to modify all high-entropy tokens.
+For example, the vast low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments.
 To address this problem, Lee et al.
-(2023) proposed an entropy-based selective watermark detection method named SWEET.
-Specifically, they only detect tokens that exceed a certain entropy threshold, thus excluding tokens that are difficult to be modified by the watermark generator.
-However, this method also has some problems.
-First, their method requires a manually set entropy threshold, which needs to be determined by analyzing samples of code written by humans.
+(2023) proposed a selective watermark detection method named SWEET.
+Specifically, they only detect tokens that exceed a certain entropy threshold.
+However, this method has some problems.
+First, their method requires a manually-set entropy threshold by analyzing samples of code written by humans.
 At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
-To address the aforementioned problems on low-entropy watermark detection, we introduced an Entropy-based Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text.
-Specifically, we argue that the importance of a token during detection should be proportional to its entropy.
+To address the aforementioned problems on low-entropy watermark detection without changing the watermark generation process, we introduced an Entropy-based Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text.
+Specifically, we argue that the influence of a token during detection should be proportional to its entropy.
 For example, the low-entropy tokens should have a limited impact on the detection result.
-Therefore, we assign an importance weight to each generated token based on its entropy, which is continuous rather than binary as in SWEET.
+Therefore, we assign an influence weight to each generated token based on its entropy, which is continuous rather than binary as in SWEET.
 Then, inside the detector, the watermark level of a given text will be the weighted sum of all the watermarked tokens.
@@
 The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations.
-Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy level.
+Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy levels.
 There is no need to switch between different detectors facing various kinds of input texts.
-In the experiment, we evaluated the detection accuracy of our proposed methods in various low-entropy generation tasks, including code generation.
-We also test the versatility of our method in text of high entropy level, such as document-to-text generation.
+In the experiment, we evaluated the detection accuracy of EWD in low-entropy code generation task.
+We also test the versatility of our method in text of high entropy levels, such as document-to-text generation.
 Overall, our method outperformed baselinse methods including KGW and SWEET in detection accuracy.
+Besides, we included evaluation of our proposed EWD against watermark-removal attacks.
 In summary, the contributions of our work are summarized as follows:
   • We proposed an entropy-based watermark detection method called EWD.
-  • We provide a theoretical analysis of the detection accuracy in our method in a low-entropy setting.
-  • Through experiments, we show that our method not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
+  • We provide a theoretical analysis of the detection accuracy in a low-entropy setting.
+  • We showed that EWD not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
 2 Related Work
-Watermark message can be embedded into existing machine-generated text by performing a series of modifications, such as lexical substitutions Munyer and Zhong (2023); Topkara et al.
+Watermark message can be embedded into existing text by performing a series of modifications, such as lexical substitutions Munyer and Zhong (2023); Topkara et al.
 (2006); Yang et al.
@@
 (2006) used WordNetFellbaum (1998) as his dictionary while Munyer and Zhong (2023) utilized a pretrained Word2Vec model to select the alternative words.
-To ensure that the repalced words would cause minimal changes to the text semantics, context-aware lexical substitution was introduced.
-For example, Yang et al.
+To ensure that the repalced words would cause minimal changes to the text semantics, Yang et al.
 (2022) proposed a BERT-based infill model to generate replacement with regard to the sentence meaning.
@@
 (2023a) fine-tuned the infill model to generate consistent word distribution given several corrupted texts of the same origin.
-However, the level of modification to an existing text is often limited due to the need to protect text quality, so the capacity and effectiveness of lexical-based watermarks are largely restricted.
+However, the level of modification to an existing text is often limited to protect text quality, thus restricting the capacity and effectiveness of lexical-based watermarks.
 Watermark embedding can also be conducted during the LLM generation phase, either by modifying the token sampling or model logits.
@@
 (2023) introduced a technique that incorporates watermark message by predefining a random number sequence for sampling the tokens.
-Detection process is to check the correlation of the input text with random number sequences.
 To enhance resistance to text modifications, Kuditipudi et al.
-(2023) utilized the Levenshtein distance in matching the text with number sequences.
+(2023) utilized the Levenshtein distance in matching the text and numbers.
 For the latter, the most representative one that modifies the model logits is proposed by Kirchenbauer et al.
 (2023a) based on previous tokens during text generation.
-Specifically, the model vocabulary will be separated into green and red lists during each generation step.
-And the logit value of every green token will be enhanced by a positive constant.
 There are a number of works proposed on top of this algorithm with enhanced performance in payload Yoo et al.
@@
 For example, Yoo et al.
-(2023b) extends the method to a multi-bit setting to allow the embedding of longer messages.
-Wang et al.
+(2023b) extends the method to a multi-bit setting.Wang et al.
 (2023) proposed to utilize a proxy LLM for multi-bit message embedding.
 To increase robustness against editing attacks, Zhao et al.
-(2023) proposed to set a universal green and red lists during the text generation process.
+(2023) proposed to set a universal green list during the generation process.
 Liu et al.
-(2023b) proposed a different approach which utilizes the semantic meaning of previous generated text to determine the green and red lists.
+(2023b) proposed a different approach which utilizes the semantic meaning of previous generated text to determine the green lists.
 Also, to enhance the unforgeability of text watermark, Liu et al.
 (2023a) introduced a neural network for text generation and detection.
-In this work, we focus more on the detection of low-entropy texts.
+In this work, we focus only on the detection of low-entropy texts.
 Lee et al.
@@
 3.1 LLM
-Here we introduce the necessary concepts used in this paper.
+Here we introduce the necessary concepts in this paper.
 A large language model (LLM) takes a prompt as input and outputs subsequent tokens as the corresponding response.
@@
 The LLM would generate a distribution logits $P_{LLM}(x^{prompt},T_{:(l-1)})$, consisting of $P_{LLM}(v|x^{prompt},T_{:(l-1)})$ over each token $v$ of its model vocabulary based on the joint input.
-The token sampling is based on the logits and specified sampling strategies, such as greedy sampling.
 3.2 Text Watermarking
@@
 4 Method
-In this section, we will first explain the motivation by analysing the watermark effect on low-entropy tokens and previously-proposed detection methods, then provide a detailed introduction to the proposed EWD algorithm.
+In this section, we will first explain the motivation by analysing the watermark effect on low-entropy tokens and traditional detection methods, then provide a detailed introduction to the proposed EWD.
 4.1 Motivation
@@
 The distribution level of the logits, represented by the entropy, matters significantly during this process.
-When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this watermark embedding process would very likely produce a green-list (watermarked) token.
+When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this logits modification would very likely produce a green-list (watermarked) token.
 An average-high-entropy text would then consist of many watermarked tokens and output a high z-score according to Eq.1.
 However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would produce a watermarked token with a much lower and constant probability.
-As a result, the number of watermarked tokens and the consequence z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1.
+As a result, the number of watermarked tokens and the consequence detection z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1.
+This is because to the detector, tokens are of the same influence, no matter how difficult it is to modify them.
 4.2 Proposed Method
 As discussed in the previous sections, it is crucial for the detector to consider each token’s entropy.
-And we take a step further than SWEET: assigning an entropy-based weight to each token, representing their significance in the detection process.
-In figure 1, we can observe that a high-entropy token has a larger probability of being watermarked or is easier to be modified and watermarked than a low-entropy token.
-Since a low-entropy token is hard to modify, then whether it is watermarked or not should have a limited impact on the overall watermark level.
-Specifically, we argue that when an easy-to-watermark token is not watermarked, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-modify token.
-That is to say, The impact of a non-watermarked token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its likelihood to be modified, and same for the impact of a watermarked token to yielding a positive detection result:
-$I(Detect(T)=\textbf{1}(t\in G)|t)\propto Entropy(t).$ (3)
-When reflected in the real world, this impact is represented by the token weight.
-Token with a larger spike entropy, meaning that it is easier to be modified, will be given a larger weight.
+And we took a step further than SWEET: assigning an entropy-based weight to each token, representing their influence in the detection process.
+In figure 1, we can observe that a high-entropy token is easier to be modified and thus has a larger probability of being watermarked than a low-entropy token.
+Since a low-entropy token is hard to modify, then whether it is watermarked or not should have a limited influence on the overall watermark level.
+Specifically, we argued that when an easy-to-modify token is not watermarked, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-modify token.
+That is to say, The influence of a non-watermarked token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its likelihood to be modified, and same for the influence of a watermarked token to yielding a positive detection result:
+$I(Detect(T)=\textbf{1}(t\in G)|t)\propto SE(t).$ (3)
+When reflected in the real world, this influence index is represented by the token weight.
+Token with a larger spike entropy, meaning that it is easier to be modified, will be given a larger influence weight.
 On the contrary, a low-entropy token will be given a much lower weight close to 0.
@@
 We first acquire the text entropy by computing the model logits of each token.
-The next step is to determine the importance weight of each token, the output of a designed function with their entropy values as inputs.
+The next step is to determine the influence weight of each token, the output of a designed function with their entropy values as inputs.
 Due to the narrow range of entropy values, our goal here is to create a bigger gap between tokens of high entropy and low entropy.
@@
 Here, we first normalize the text entropy.
-Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights:
-$W=Linear(Norm(SE)).$ (5)
-After obtaining the weights for each token, we then apply the standard detection procedure by Kirchenbauer et al.
+Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights.
+After obtaining the entropy-based weights for each token, we then apply the standard detection procedure by Kirchenbauer et al.
 (2023a) to determine the group of watermarked tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens.
@@
 The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
-In real-world low-entropy scenarios, such as code generation, there are a large number of hard-to-modify tokens.
-When detecting watermarked texts with our proposed EWD, the vast majority of non-watermarked low-entropy tokens will not prevent the detector from yielding a positive conclusion.
-Meanwhile, facing non-watermarked texts, any non-watermarked high-entropy token will strongly suggest the detector to make a negative output.
+In real-world low-entropy scenarios, With our EWD, the vast majority of non-watermarked low-entropy tokens will not prevent the detector from yielding a positive conclusion.
+Meanwhile, facing human-written texts, any non-watermarked high-entropy token will strongly suggest the detector to make a negative output.
 5 Theoretical Analysis
-In this section, we theoretically analyzed the Type-I and Type-II error of our proposed detection method.
-Specifically, we compared EWD with two previous methods, KGWKirchenbauer et al.
-(2023a) and SWEETLee et al.
-(2023).
+In this section, we theoretically analyzed the Type-I and Type-II error of our proposed detection method. Specifically, we compared EWD with two previous detection methods, KGW and SWEET.
 To do so, we adopted Kirchenbauer et al.
@@
 Each token in both watermarked and non-watermarked texts can either be green or not, so we can model it as a binomial distribution.
-Therefore, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes.
+While Zhao et al.
+(2023) set window size in the watermark generator as zero and exhibited major token dependence, $i.e.$, the same tokens in the text are of same watermark status (either all green or all red), in our case, where a non-zero window size is adopted, the token dependence is significantly weakened.
+For example, the same token with different prefix would exhibit different watermark status.
+Therefore, to simplify and facilitate the process of theoretical analysis, we made the reasonable assumption that the watermark status of each token is independent.
+Hence, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes.
 Finally, we calculate and compare the probability of the sum surpassing or falling behind the detection threshold $|\tilde{s}|_{G}$, which is the $|s|_{G}$ value in each detection function with a fixed $z$.
@@
 When $z$ is set to 2, the probabilities for $|s|_{G}$ under each method to surpass its detection threshold $|\tilde{s}|_{G}$ is identically 2.28%.
-Table 1: Main results on different datasets using KGW, SWEET and our method for detection.
+Table 1: Main results on different datasets using KGW, SWEET and our EWD for detection.
 Code Detection (Low-entropy Scenario)
@@
 Type-II error measures the likelihood a machine-generated text is incorrectly identified as human-written.
-For the purposes of analysis, we assumed that a sequence generated by the watermark algorithm has a length of $T=200$ tokens and that the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
+For analysis, we assumed that a sequence generated by the watermark algorithm has a length of $T=200$ tokens and the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
 Originally, Kirchenbauer et al.
-(2023a) sampled over 500 watermark-generated new scripts and reported an average spike entropy of 0.807.
+(2023a) sampled over 500 watermark-generated news scripts and reported an average spike entropy of 0.807.
 In this section, we focused specifically on cases with a much lower average spike entropy.
@@
 When the average entropy $SE^{*}$ given by the simulated power law distribution is 0.607, we approximated $|s|_{G}$ by $\mathcal{N}(106.9,7.1)$.
-If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 114) for detection, the false negative rate exceeds 84.6%, or in other words, the detection rate of watermarked text is less than 15.4%.
+If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 114) for detection, the false negative rate exceeds 84.6%.
 This means that KGW is not good at detecting the watermarked texts in cases of low entropy.
 SWEET.
-Following the detection scheme in SWEET, we set the entropy threshold at 0.695 and excluded all tokens below this threshold during detection.
+Following SWEET, we set the entropy threshold at 0.695 and excluded all tokens below this threshold during detection.
 The average spike entropy $SE^{*}$ and remaining text length $\tilde{T}$ then changed to 0.82 and 24 respectively.
 Using the equations from KGW, we can get the approximation as $\mathcal{N}(17.5,2.2)$.
-If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 17.2) for detection, the false negative rate exceeds 42%, or in other words, the detection rate of watermarked text is around 58%.
+If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 17.2) for detection, the false negative rate exceeds 42%.
 EWD.
-We assigned higher weights to tokens with high spike entropy in order to achieve a better detection rate in cases of low average spike entropy.
-Note that in experiments, we use several weight functions, and here, for analysis purposes, we used a linear weight function $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant.
+We assigned higher influence weights to tokens with high spike entropy in order to achieve a better detection rate in cases of low average spike entropy.
+Note that in experiments, we use several weight functions, and here, for analysis purposes, we used a linear weight function for each token $k$: $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant.
 Based on Eq.5, the statistical $\mu$ and $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ and $\sum_{k}{W(k)}^{2}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$.
 Therefore, we can approximate the $|s|_{G}$ detected by EWD with $\mathcal{N}(5.79,0.56)$.
-If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.3%, or in other words, the detection rate of watermarked text is around 66.7% (Proof and details in Appendix A).
-Overall, Our method has a type-II error of 33.3%, much lower than KGW and SWEET.
-Empirical experiments also demonstrate similar results.
+If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.3% (Proof and details in Appendix A).
+Overall, Our method has a type-II error much lower than KGW and SWEET.
 6 Experiments
@@
 Each data point can correspond to the illustrated function on the left figure by looking at its shape and x-axis value.
-Tasks and Datasets. First of all, we used the KGW Kirchenbauer et al. (2023a) for watermark generation in all the following tasks.
+Tasks and Datasets. First of all, we used the KGW Kirchenbauer et al. (2023a) to embed watermark in all the following tasks.
 To evaluate the detection performance of our proposed method, we targeted at a commonly-seen low-entropy scenario, code generation.
@@
 During evaluation, we restricted the minimal length of both samples to be 15 tokens.
-Besides code generation, we fine-tuned a T5-small Raffel et al.
+Besides, we fine-tuned a T5-small Raffel et al.
 (2023) model for data-to-text generation.
@@
 The dataset consists of NBA game records, presented as tables, and human-written descriptions.
-We used the table information as input and asked the fine-tuned T5 to generate descriptive texts by binomial sampling and beam-search.
+We used the table information as input and generate descriptive texts.
 For each input prompt, the models are expected to generate 200 ± 5 tokens.
@@
 Baselines and Evaluation Metrics.
-We selected two detection methods as the baselines to compare with.
-The first is KGW Kirchenbauer et al.
-(2023a), which is the original detector.
-The second is SWEET Lee et al.
+We selected two detection methods as baselines.
+The first is KGW detector Kirchenbauer et al.
+(2023a).
+The second is SWEET’s detector Lee et al.
 (2023), which excludes low-entropy tokens during detection.
-We follow the instructions in SWEET to select entropy threshold, namely from the median, mean, and 3rd quartile (Q3) values of the entropy computed from human texts.
+We follow the instructions in SWEET to select entropy threshold from the median, mean, and 3rd quartile (Q3) values of the entropy computed from human texts.
 For evaluation, we set the false positive rate (FPR) under 5%, following Zhao et al.
@@
 Hyper-parameters.
-For watermarked generation, we set $\gamma$ and $\delta$ to 0.5 and 2 respectively.
-In text generation using binomial sampling, we set the sampling temperature to 0.7.
-In text generation using beam-search, we set the number of beams to 8.
-In order to avoid excessive repetition of text, we set "no_repeated_ngrams" to 16.
+For watermarked algorithm, we set $\gamma$ and $\delta$ to 0.5 and 2 respectively.
+Binomial sampling set the sampling temperature to 0.7.
+Beam-search set the number of beams to 8.
+To avoid excessive repetition of text, we set "no_repeated_ngrams" to 16.
 6.2 Main Results
@@
 (2023)’s.
-For the Rotowire dataset, our generated output can achieve 7.24 in text perplexity (PPL).
-For the C4 dataset, our generated output can achieve 3.25 in PPL.
-PPL is calculated using OPT-6.7B Zhang et al.
+For the Rotowire and C4 dataset, our generated output can achieve 7.24 and 3.25 in text perplexity (PPL) using OPT-6.7B Zhang et al.
 (2022).
@@
 Impact of Entropy-based Weights.
-In figure 2, we give a few visualizations to show the reason why our method surpassed other baselines in detection performance.
+In figure 2, we gave a few visualizations to show the reason why EWD surpassed other baselines in detection performance.
 Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods.
-Our detection methods would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts.
+Our EWD method would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts.
 The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset.
 By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts.
-In figure 2(b), we plotted the relationship of entropy-based weight and probability of being watermarked.
+In figure 2(b), we plotted the relationship of a token’s entropy-based weight and probability of that token being watermarked.
 In watermarked text, the probability is proportional to the token weight.
@@
 Specifically, our detector would also compute weight given the token entropy before calculating the final z-score.
-According to table 2, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our detector respectively.
+According to table 2, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our EWD respectively.
 In spite of being double the value of KGW, our detector remains highly efficient, and the difference with SWEET is virtually insignificant.
+Performance against Attacks.
+Table 3: Detection performance of back-translated watermarked texts using different detection methods.
+C4 w/ Back-translation
+Methods 1%FPR 5%FPR
+TPR F1 TPR F1
+KGW 0.894 0.940 0.950 0.952
+SWEET 0.890 0.938 0.930 0.940
+EWD 0.900 0.943 0.942 0.947
+Watermarked texts usually would be edited, which would cause detection performance drop.
+Here we also evaluated the baseline methods and our EWD under this setting by utilizing back-translation as the attack to remove watermark.
+Specifically, we firstly generated watermarked texts using the C4 dataset, then translated them from English to French, and later back to English again for detection.
+In table 3, we can observe that our EWD would achieve the best detection accuracy under 1% FPR and similar results with KGW under 5% FPR.
 7 Conclusion

### v3

@@
 Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved.
-In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token during watermark detection should be adjusted according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods.
-Specifically, we proposed an Entropy-based Watermark Detection (EWD) that gives higher-entropy tokens higher weights during watermark detection, so as to better reflect the degree of watermarking.
+In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token during watermark detection should be customized according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods.
+Specifically, we proposed an Entropy-based Text Watermark Detection (EWD) that gives higher-entropy tokens higher influence weights during watermark detection, so as to better reflect the degree of watermarking.
 Furthermore, the proposed detection process is training-free and fully automated.
 In the experiment, we found that our method can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions.
-Our code and data will be available online.
+Our code and data is available on https://github.com/luyijian3/EWD.
+Additionally, our algorithm could be accessed through MarkLLMPan et al.
+(2024).
 An Entropy-based Text Watermarking Detection Method
-Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1 1The Chinese University of Hong Kong 2Tsinghua University
+Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1 1The Chinese University of Hong Kong 2Tsinghua University luyijian@link.cuhk.edu.hk, liuaw20@mails.tsinghua.edu.cn king@cse.cuhk.edu.hk
 1 Introduction
 Figure 1: Watermarking algorithm has limited effect on a low-entropy token as compared to a high-entropy token.
-Therefore, a low-entropy token has a much smaller probability of being watermarked.
-In a code generation scenario, most of the tokens are low-entropy, resulting in less watermarked tokens and a small detection z-score.
+Therefore, it is almost random for a low-entropy token to be green.
+In a code generation scenario, most of the tokens are low-entropy, resulting in less green tokens and a small detection z-score.
 The advancements in LLMs have enabled it to generate high-quality outputs indistinguishable from humans, and also achieve better performance on various real-world generation tasks.
@@
 The current LLM watermarking technology is usually to embed watermark information in the process of generating tokens during LLM inference.
-For example, Kirchenbauer et al.
+For example, KGW Kirchenbauer et al.
 (2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list.
-By modifying the token distributions, it leads the generated text to contain mostly watermarked (green) tokens.
-In the watermark detection process, the degree of watermark is usually judged by the number of watermarked tokens in the text.
+By modifying the token distributions, it imposes a preference on the green list to the model and leads the model to sample most tokens from the green list.
+In the detection process, the degree of watermark is usually judged by the number of green tokens in the text.
 Although current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved.
-The reason behind this is that low-entropy tokens usually have stronger determinism, are more difficult to modify in the watermarking process, and are easier to affect the quality of the text when forcefully modified Tu et al.
+The reason behind this is that low-entropy tokens usually have stronger determinism, and the modified logits can hardly affect the probability of sampling the tokens from the green list.
+At the same time, if logits is forcefully modified, the quality of the watermarked text is easier to degrade Tu et al.
 (2023).
-In figure 1, we illustrate the watermark ratio in tokens of three different entropy levels.
-A token with higher entropy will have a significantly larger probability to be modified and watermarked.
-On the contrary, it is rare to see a green token in low-entropy tokens.
-When the generated text contains a large number of low-entropy tokens, tokens that can hardly be modified by the watermarking algorithm, the detection results of the watermark detector are more easily affected by the states of these unmodified low-entropy tokens even when the watermark generator managed to modify all high-entropy tokens.
+In figure 1, we illustrate the green-list ratio in tokens of three different entropy levels.
+A token with higher entropy will have a significantly larger probability to be sampled from a green list.
+On the contrary, it is almost random to see a green token in low-entropy tokens if the green list size is half of the model vocabulary size.
+When the generated text contains a large number of low-entropy tokens, tokens that can hardly be affected by the modified logits, the detection results of the watermark detector are more easily affected by the states (colors) of these low-entropy tokens even when the watermark generator managed to sample all high-entropy tokens from the green list.
 For example, the vast low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments.
@@
 At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
-To address the aforementioned problems on low-entropy watermark detection without changing the watermark generation process, we introduced an Entropy-based Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text.
+To address the aforementioned problems on low-entropy watermark detection without changing the watermark generation process, we introduced an Entropy-based Text Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text.
 Specifically, we argue that the influence of a token during detection should be proportional to its entropy.
 For example, the low-entropy tokens should have a limited impact on the detection result.
-Therefore, we assign an influence weight to each generated token based on its entropy, which is continuous rather than binary as in SWEET.
-Then, inside the detector, the watermark level of a given text will be the weighted sum of all the watermarked tokens.
-By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm failed to watermark a low-entropy token, it will make minimal impact to the detected watermark level.
+Furthermore, in order to fully reflect this positive relation between token entropy and detection influence, we utilized a monotonically-increasing and continuous function to generate influence weights from token entropy.
+Then, inside the detector, the watermark level of a given text will be the weighted sum of all the green tokens.
+By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm can hardly affect a low-entropy token, the state of that token will make minimal impact to the detected watermark result.
 The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations.
@@
 In summary, the contributions of our work are summarized as follows:
-  • We proposed an entropy-based watermark detection method called EWD.
+  • We proposed an entropy-based text watermark detection method called EWD.
   • We provide a theoretical analysis of the detection accuracy in a low-entropy setting.
@@
 (2024), specifically the ones concerning Logits Generation.
-The general watermarking scheme used in this paper is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$.
+The general watermarking scheme used in this paper is called KGW Kirchenbauer et al.
+(2023a), which is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$.
 The modification on the original logits by adding a small value to the logits of green-list tokens will result in LLM’s preference on those tokens in the output text.
-Thus such modification could likely produce a watermarked (green) token.
+Thus such modification could likely produce a green token.
 The detection algorithm is to calculate the watermark level using the following equation
 $z=(|s|_{G}-\gamma T)/\sqrt{\gamma(1-\gamma)T},$ (1)
-where $|s|_{G}$ is the number of watermarked tokens and $\gamma$ is the green-list ratio in the model vocabulary.
+where $|s|_{G}$ is the number of green tokens and $\gamma$ is the green-list ratio in the model vocabulary.
 The algorithm would output 1 if the value is significantly greater than a given threshold, and 0 otherwise.
@@
 (2023a),
-$SE(k)=\sum_{k}{\frac{p_{k}}{1+zp_{k}}},$ (2)
-where $p_{k}$ is the probability vector of the token $k$ and $z$ is a scalar.
+$SE(k)=\sum_{v}{\frac{p_{v}}{1+zp_{v}}},$ (2)
+where $p$ is the logits when sampling the token $k$ and $z$ is a scalar.
 The entropy value achieves minimal if the logits concentrates on a single token and maximal given a uniform distribution.The watermark generation process is more effective on a high-entropy token where the logits distribution is nearly uniform, so that the modification could produce a green-list token easily.
 On the contrary, a low-entropy token is almost deterministic and hard to be modified.
-Therefore, the watermarked ratio for a low-entropy token is much lower, resulting in fewer watermarked tokens in the output.
+Therefore, the green-list ratio for a low-entropy token is much lower, resulting in fewer green tokens in the output.
 A low-entropy scenario refers to the LLM-generated output consisting of a great number of low-entropy tokens.
@@
 The distribution level of the logits, represented by the entropy, matters significantly during this process.
-When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this logits modification would very likely produce a green-list (watermarked) token.
-An average-high-entropy text would then consist of many watermarked tokens and output a high z-score according to Eq.1.
-However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would produce a watermarked token with a much lower and constant probability.
-As a result, the number of watermarked tokens and the consequence detection z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1.
-This is because to the detector, tokens are of the same influence, no matter how difficult it is to modify them.
+When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this logits modification would very likely produce a green-list token.
+An average-high-entropy text would then consist of many green tokens and output a high z-score according to Eq.1.
+However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would hardly affect the token sampling and produce a green token with a much lower probability close to the green list ratio $\gamma$.
+As a result, the number of green tokens and the consequence detection z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1.
+This is because to the detector, tokens are of the same influence, no matter how difficult it is for the logits modification to affect them.
 4.2 Proposed Method
-As discussed in the previous sections, it is crucial for the detector to consider each token’s entropy.
-And we took a step further than SWEET: assigning an entropy-based weight to each token, representing their influence in the detection process.
-In figure 1, we can observe that a high-entropy token is easier to be modified and thus has a larger probability of being watermarked than a low-entropy token.
-Since a low-entropy token is hard to modify, then whether it is watermarked or not should have a limited influence on the overall watermark level.
-Specifically, we argued that when an easy-to-modify token is not watermarked, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-modify token.
-That is to say, The influence of a non-watermarked token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its likelihood to be modified, and same for the influence of a watermarked token to yielding a positive detection result:
+In figure 1, we can observe that a high-entropy token is easier to be affected by the modified logits, and has a larger probability of being sampled from the green list.
+Since a low-entropy token is hard to affect, then whether it is sampled from the green list or not should have a limited influence on the overall watermark level.
+Specifically, we argued that when an easy-to-affect (high-entropy) token is not green, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-affect (low-entropy) token.
+That is to say, The influence of a red token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its entropy, and same for the influence of a green token to yielding a positive detection result:
 $I(Detect(T)=\textbf{1}(t\in G)|t)\propto SE(t).$ (3)
 When reflected in the real world, this influence index is represented by the token weight.
-Token with a larger spike entropy, meaning that it is easier to be modified, will be given a larger influence weight.
+Token with a larger spike entropy, meaning that it is easier to be affected, will be given a larger influence weight.
 On the contrary, a low-entropy token will be given a much lower weight close to 0.
-As the sum of all watermarked token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a non-watermarked low-entropy token than a non-watermarked high-entropy token, and a larger gain when observing a watermarked high-entropy token than a watermarked low-entropy token.
+As the sum of all green token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a red low-entropy token than a red high-entropy token, and a larger gain when observing a green high-entropy token than a green low-entropy token.
 Algorithm 1 Entropy-based Watermark Detection
@@
 else return 0, $i.e.,$Input text is not watermarked"
+Now we are aware of the positive influence-entropy relation as suggest Eq.3.
+We decided to take a step further than SWEET: customizing influence weight according to token entropy to fully reflect the influence-entropy relation.
+While in SWEET, the weight of 1 for above-threshold tokens and the weight of 0 for below-threshold tokens partially reflect the relation, there is not reflection inside the entropy range of above-threshold tokens.
+Therefore, we utilized a monotonically-increasing and continuous function $f$ to generate weights from token entropy.
+The weight assigning function can be illustrated as follows:
+$W(t)=f(SE(t)-C_{0}).$ (5)
 Pseudocode of our proposed detection method is provided in Algorithm 1.
@@
 After obtaining the entropy-based weights for each token, we then apply the standard detection procedure by Kirchenbauer et al.
-(2023a) to determine the group of watermarked tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens.
-Finally, we sum up the weights of the watermarked tokens and calculate the z-statistic using Eq.4 given the sum and weight vector of all detected tokens.
+(2023a) to determine the group of green tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens.
+Finally, we sum up the weights of the green tokens and calculate the z-statistic using Eq.4 given $|s|_{G}$ and weights of all detected tokens.
 The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
-In real-world low-entropy scenarios, With our EWD, the vast majority of non-watermarked low-entropy tokens will not prevent the detector from yielding a positive conclusion.
-Meanwhile, facing human-written texts, any non-watermarked high-entropy token will strongly suggest the detector to make a negative output.
+In real-world low-entropy scenarios, With our EWD, the vast majority of low-entropy red tokens will not prevent the detector from yielding a positive conclusion.
+Meanwhile, facing human-written texts, any high-entropy red token will strongly suggest the detector to make a negative output.
 5 Theoretical Analysis
@@
 While Zhao et al.
-(2023) set window size in the watermark generator as zero and exhibited major token dependence, $i.e.$, the same tokens in the text are of same watermark status (either all green or all red), in our case, where a non-zero window size is adopted, the token dependence is significantly weakened.
-For example, the same token with different prefix would exhibit different watermark status.
-Therefore, to simplify and facilitate the process of theoretical analysis, we made the reasonable assumption that the watermark status of each token is independent.
+(2023) set window size in the watermark generator as zero (one universal green list for all tokens) and exhibited major token dependence, $i.e.$, the same tokens in the text are of the same color (either all sampled from the green list or all from the red list), in our case, where a non-zero window size is adopted, the token dependence is significantly weakened.
+For example, the same token with different prefix would exhibit different sampling result.
+Therefore, to simplify and facilitate the process of theoretical analysis, we made the reasonable assumption that the sampling of each token is independent.
 Hence, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes.
@@
 (2023) that only considers high-entropy tokens, such distribution is modelled as $\mathcal{N}(\gamma\tilde{T},\gamma(1-\gamma)\tilde{T})$, where $\tilde{T}$ is the text length excluding low-entropy tokens.
-Our method EWD added weights to tokens based on entropy, leading to a weighted binomial distribution.
+Our method EWD assigned weights to tokens based on entropy, leading to a weighted binomial distribution.
 We could also approximate it to a normal distribution with mean $\mu=\gamma\sum_{i=m}^{|T|-1}{W}_{i}$ and variance $\sigma^{2}=\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}$.
@@
 TPR F1 TPR F1 F1 TPR F1 TPR F1 F1
-KGW 0.383 0.551 0.519 0.663 0.776 0.113 0.201 0.327 0.477 0.744
-SWEET 0.662 0.793 0.744 0.835 0.897 0.482 0.647 0.692 0.795 0.876
-EWD 0.684 0.809 0.850 0.897 0.900 0.615 0.757 0.748 0.833 0.888
+KGW 0.331 0.494 0.414 0.567 0.772 0.115 0.205 0.348 0.498 0.744
+SWEET 0.455 0.622 0.667 0.778 0.842 0.409 0.577 0.650 0.765 0.865
+EWD 0.466 0.633 0.692 0.797 0.859 0.567 0.720 0.744 0.830 0.878
 Rotowire (High-entropy Scenario)
@@
 TPR F1 TPR F1 F1 TPR F1 TPR F1 F1
-KGW 0.986 0.988 0.993 0.972 0.990 1.000 0.996 1.000 0.976 0.999
-SWEET 0.986 0.988 0.993 0.972 0.990 0.997 0.994 1.000 0.976 0.996
-EWD 0.984 0.987 0.997 0.974 0.988 0.999 0.995 1.000 0.976 0.997
+KGW 1.000 0.995 1.000 0.976 1.000 1.000 0.995 1.000 0.976 1.000
+SWEET 1.000 0.995 1.000 0.976 1.000 1.000 0.995 1.000 0.976 1.000
+EWD 1.000 0.995 1.000 0.976 1.000 1.000 0.995 1.000 0.976 1.000
 C4 (High-entropy Scenario)
@@
 TPR F1 TPR F1 F1 TPR F1 TPR F1 F1
-KGW 0.998 0.995 0.999 0.976 0.998 0.998 0.995 0.999 0.976 0.998
-SWEET 0.995 0.993 0.997 0.975 0.993 0.995 0.993 0.997 0.975 0.993
-EWD 0.996 0.993 0.997 0.975 0.996 0.996 0.993 0.997 0.975 0.996
+KGW 0.985 0.990 1.000 0.980 0.992 0.995 0.993 1.000 0.980 0.997
+SWEET 1.000 0.997 1.000 0.978 1.000 0.995 0.993 1.000 0.976 0.997
+EWD 1.000 0.997 1.000 0.978 1.000 0.995 0.993 1.000 0.976 0.997
 5.2 Type-II Error
@@
 In this section, we focused specifically on cases with a much lower average spike entropy.
-Based on our data, we approximated the distribution of spike entropy with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$.
+Based on our data, we approximated the distribution of spike entropy in low-entropy watermarked texts with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$.
 Under this distribution, the average spike entropy is 0.607.
 The following probability from Kirchenbauer et al.
-(2023a) showed a lower bound of the probability of a token $k$ being watermarked by a watermark generator:
+(2023a) showed a lower bound of the probability of a token $k$ being sampled from the green list:
 $\mathbb{P}[k\in G]\geq\frac{\gamma\alpha}{1+(\alpha-1)\gamma}SE(k),$ (6)
@@
 Figure 2: The left figure shows the z-scores of watermarked and human texts in the Rotowire, HumanEval and MBPP datasets respectively, each being detected with 3 different methods.
-The right figure shows the relationship between token weights and the probability of being watermarked in both watermarked and human texts.
+The right figure shows the relationship between token weights and the probability of being sampled from the green list in both watermarked and human texts.
 Figure 3: We utilized a variety of weight functions and measured their performance in the code detection datasets.
 The left figure is a visualization of all studied functions, with normalized spike entropy as input.
-The right figure shows each function’s detection F1 under 5% FPR with comparison to the SWEET baseline.
+The right figure shows each function’s detection F1 under 1% FPR with comparison to the SWEET baseline.
 Each data point can correspond to the illustrated function on the left figure by looking at its shape and x-axis value.
@@
 (2022).
-Table 1 shows that for the low-entropy scenario, code detection, our detector would achieve the best detection accuracy against other baselines.
-Specifically, for HumanEval and MBPP respectively, our method could achieve a 10.6% and 5.6% improvement than SWEET in true positive rate while maintaining a low-than 5% false positive rate.
-Table 1 shows that for the high-entropy scenarios using Rotowire and C4 datasets, our method is better than SWEET and is very similar to the original KGW method, with very minor differences.
+Table 1 shows that for the low-entropy scenario, watermarked code detection, our detector would achieve the best detection accuracy against other baselines.
+Specifically, for HumanEval and MBPP respectively, our method could achieve a 3.5% and 9.4% improvement than SWEET in true positive rate while maintaining a low-than 5% false positive rate.
+Table 1 shows that for the high-entropy scenarios using Rotowire and C4 datasets, our method is overall very similar to other baselines, sometimes even better than them.
 6.3 Analysis
@@
 By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts.
-In figure 2(b), we plotted the relationship of a token’s entropy-based weight and probability of that token being watermarked.
+In figure 2(b), we plotted the relationship of a token’s entropy-based weight and probability of that token being sampled from the green list.
 In watermarked text, the probability is proportional to the token weight.
-That is to say, the tokens assigned with a large weight are very likely to be watermarked and contribute to a larger z-score.
+That is to say, the tokens assigned with a large weight are very likely to be sampled from the green list and contribute to a larger z-score.
 On the contrary, in human-written texts, there doesn’t exist such a relationship, explaining why the z-scores of human text remain low.
@@
 In the experiments, we utilized a linear function to gain the entropy-based weights for each token.
-We also studied a variety of monotonic increasing functions as shown in figure 3.
+We also studied a variety of monotonic increasing functions as shown in figure 3, which are applied in Eq.5 as $f$.
 Specifically, each function’s gradient is different.
@@
 Similarly for our custom Exponential functions, the gradient starts off low and gradually increases, meaning that tokens with relatively low entropy will have smaller differences in weight, while tokens with relatively high entropy will exhibit larger differences.
-We use function strength to represent how far the function deviates from being linear, for example, "Exp 10" deviates from linear more than "Exp 8".
+We use function strength to represent how far the function deviates from being linear, for example, "Sigmoid 10" deviates from linear more than "Sigmoid 8".
 In the experiments with those functions and two code generation datasets, we can observe that the total 8 listed functions could surpass the SWEET baseline in terms of detection F1.
@@
 SWEET 9.489 0.0830
-Ours 0.0831
+EWD 0.0831
 Compared to the KGW detection, both SWEET and our proposed EWD would demand extra time to compute token entropy.
@@
 Table 3: Detection performance of back-translated watermarked texts using different detection methods.
-C4 w/ Back-translation
+Detection w/ Back-translation
 Methods 1%FPR 5%FPR
@@
 In table 3, we can observe that our EWD would achieve the best detection accuracy under 1% FPR and similar results with KGW under 5% FPR.
+[…]
+Table 4: Detection performance of watermarked code using the general prompt instead of original prompts.
+Detection w/ General Prompt
+Method F1 under 5%FPR Best F1
+KGW 0.498 0.741
+SWEET 0.580 0.762
+EWD 0.600 0.764
 7 Conclusion

### v4

@@
 Abstract
-Currently, text watermarking algorithms for large language models (LLMs) can embed hidden features to texts generated by LLMs to facilitate subsequent detection, thus alleviating the problem of misuse of LLMs.
+Text watermarking algorithms for large language models (LLMs) can effectively identify machine-generated texts by embedding and detecting hidden features in the text.
 Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved.
-In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token during watermark detection should be customized according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods.
-Specifically, we proposed an Entropy-based Text Watermark Detection (EWD) that gives higher-entropy tokens higher influence weights during watermark detection, so as to better reflect the degree of watermarking.
+In this work, we opine that the influence of token entropy should be fully considered in the watermark detection process, $i.e.$, the weight of each token during watermark detection should be customized according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods.
+Specifically, we propose Entropy-based Text Watermarking Detection (EWD) that gives higher-entropy tokens higher influence weights during watermark detection, so as to better reflect the degree of watermarking.
 Furthermore, the proposed detection process is training-free and fully automated.
-In the experiment, we found that our method can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions.
-Our code and data is available on https://github.com/luyijian3/EWD.
+From the experiments, we demonstrate that our EWD can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions.
+Our code and data is available [^1] https://github.com/luyijian3/EWD.
 Additionally, our algorithm could be accessed through MarkLLM Pan et al.
-(2024).
+(2024) [^2] https://github.com/THU-BPM/MarkLLM.
 An Entropy-based Text Watermarking Detection Method
-Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1 1The Chinese University of Hong Kong 2Tsinghua University luyijian@link.cuhk.edu.hk, liuaw20@mails.tsinghua.edu.cn king@cse.cuhk.edu.hk
+Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1thanks: Corresponding Author.
+1The Chinese University of Hong Kong 2Tsinghua University luyijian@link.cuhk.edu.hk, liuaw20@mails.tsinghua.edu.cn dianzhi.yu@link.cuhk.edu.hk, lijj@link.cuhk.edu.hk, king@cse.cuhk.edu.hk
 1 Introduction
-Figure 1: Watermarking algorithm has limited effect on a low-entropy token as compared to a high-entropy token.
-Therefore, it is almost random for a low-entropy token to be green.
-In a code generation scenario, most of the tokens are low-entropy, resulting in less green tokens and a small detection z-score.
-The advancements in LLMs have enabled it to generate high-quality outputs indistinguishable from humans, and also achieve better performance on various real-world generation tasks.
+Figure 1: This figure shows that compared with watermarked texts with mostly high-entropy tokens, watermarked codes with mostly low-entropy tokens see significantly less green tokens, resulting in a small detection z-score.
+Furthermore, on the bottom of the figure, we demonstrate that the green token ratio in tokens decreases as their entropy decreases.
+The rapid advancements in large language models (LLMs) have enabled them to generate high-quality outputs indistinguishable from humans, and also achieve better performance on various real-world generation tasks.
 However, this gives rise to the potential risk of misuse.
-Specifically, LLM-generated assignments, including essays, code, and reports, pose a threat to academic integrity.
+Specifically, LLM-generated assignments, including essays and codes, pose a threat to academic integrity.
 Misusing LLMs to generate fake news can also lead to negative societal consequences Megías et al.
 (2021).
-Therefore, algorithms capable of effectively marking and identifying machine-generated contents have become essential.
-The text watermarking algorithm for LLMs can embed hidden features that can be detected later during the text generation process of LLMs, thus effectively alleviating the problem of misuse of LLMs.
-The current LLM watermarking technology is usually to embed watermark information in the process of generating tokens during LLM inference.
-For example, KGW Kirchenbauer et al.
-(2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list.
-By modifying the token distributions, it imposes a preference on the green list to the model and leads the model to sample most tokens from the green list.
-In the detection process, the degree of watermark is usually judged by the number of green tokens in the text.
-Although current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved.
-The reason behind this is that low-entropy tokens usually have stronger determinism, and the modified logits can hardly affect the probability of sampling the tokens from the green list.
-At the same time, if logits is forcefully modified, the quality of the watermarked text is easier to degrade Tu et al.
+Therefore, algorithms capable of effectively identifying machine-generated contents have become essential.
+Text watermarking algorithms can effectively alleviate the problem of LLM misuse by embedding and detecting hidden features in LLM-generated contents.
+For example, the watermarking algorithm proposed by Kirchenbauer et al.
+(2023a) (KGW) divides the LLM’s model vocabulary into two lists, green and red, and increases the logits value of tokens in the green list by a positive constant (watermark strength).
+The modified logits will bias the LLM to sample most tokens from the green list.
+As a result, the watermarked text containing a dominant number of green tokens will be recognized by the watermark detector while human texts will not.
+However, the successful watermark generation is largely tied to the token entropy during text generation.
+Under the watermark algorithm, the probability of sampling a token from the green list is proportional to its entropy, meaning that low-entropy tokens would produce far less green tokens than high-entropy tokens, as illustrated by Figure 1.
+In that case, green tokens will occupy only half of a watermarked low-entropy text, resulting in a smaller detection score and being mistakenly labeled as human-written.
+Furthermore, to ensure the text quality of output contents, the watermark strength in the generator cannot be excessively enlarged to forcefully modify the logits and sample more tokens from the green list Tu et al.
 (2023).
-In figure 1, we illustrate the green-list ratio in tokens of three different entropy levels.
-A token with higher entropy will have a significantly larger probability to be sampled from a green list.
-On the contrary, it is almost random to see a green token in low-entropy tokens if the green list size is half of the model vocabulary size.
-When the generated text contains a large number of low-entropy tokens, tokens that can hardly be affected by the modified logits, the detection results of the watermark detector are more easily affected by the states (colors) of these low-entropy tokens even when the watermark generator managed to sample all high-entropy tokens from the green list.
-For example, the vast low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments.
-To address this problem, Lee et al.
-(2023) proposed a selective watermark detection method named SWEET.
-Specifically, they only detect tokens that exceed a certain entropy threshold.
-However, this method has some problems.
+Consequently, a crucial problem is how to improve the watermark detector so that low-entropy watermarked texts with a limited number of green tokens can be successfully identified.
+Lee et al.
+(2023) propose a selective watermark detector named SWEET to better reflect the watermark level in low-entropy texts.
+Specifically, they only detect tokens that exceed a certain entropy threshold, thus excluding tokens that are difficult to be affected by the watermark generator.
+However, this method itself has three weaknesses.
 First, their method requires a manually-set entropy threshold by analyzing samples of code written by humans.
-At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
-To address the aforementioned problems on low-entropy watermark detection without changing the watermark generation process, we introduced an Entropy-based Text Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text.
-Specifically, we argue that the influence of a token during detection should be proportional to its entropy.
-For example, the low-entropy tokens should have a limited impact on the detection result.
-Furthermore, in order to fully reflect this positive relation between token entropy and detection influence, we utilized a monotonically-increasing and continuous function to generate influence weights from token entropy.
-Then, inside the detector, the watermark level of a given text will be the weighted sum of all the green tokens.
-By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm can hardly affect a low-entropy token, the state of that token will make minimal impact to the detected watermark result.
-The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations.
-Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy levels.
-There is no need to switch between different detectors facing various kinds of input texts.
-In the experiment, we evaluated the detection accuracy of EWD in low-entropy code generation task.
-We also test the versatility of our method in text of high entropy levels, such as document-to-text generation.
-Overall, our method outperformed baselinse methods including KGW and SWEET in detection accuracy.
-Besides, we included evaluation of our proposed EWD against watermark-removal attacks.
+Second, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly vary.
+Third, the performance gain using the SWEET detector is limited and there is still significant room for further advancement.
+To better address the low-entropy watermarking detection problem than previous methods, we introduce an Entropy-based Text Watermark Detection (EWD) algorithm to fully consider the role of entropy in the detection of a given text.
+Specifically, we propose that the influence of a token during detection should be proportional to its entropy.
+Furthermore, in order to fully reflect this positive relation between token entropy and detection influence, we utilize a monotonically-increasing and continuous function to generate influence weights from token entropy.
+By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the watermark generation algorithm can hardly affect a low-entropy token, the state of that token will make minimal impact to the detection result.
+Overall, our method has three major benefits.
+First, the method saves the trouble of altering the watermark generation process.
+Second, compared to SWEET, our method is more efficient as the acquisition of token weights is an independent and training-free process, avoiding the trouble of a human dataset or any manual operations.
+Third, the proposed detection scheme can be applied to machine-generated texts of all entropy levels, not solely low-entropy texts.
+From the experiments, we demonstrate that the detection accuracy of EWD in low-entropy scenarios surpasses existing baselines.
+Furthermore, EWD exhibits versatility in detecting texts of high entropy levels, as well as robustness against the back-translation watermark-removal attack.
 In summary, the contributions of our work are summarized as follows:
-  • We proposed an entropy-based text watermark detection method called EWD.
-  • We provide a theoretical analysis of the detection accuracy in a low-entropy setting.
-  • We showed that EWD not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
+  • We propose an entropy-based text watermarking detection method called EWD.
+  • We provide a theoretical analysis of the detection accuracy in the low-entropy setting.
+  • We show that EWD not only improves the detection accuracy under the low-entropy scenario, but also provides similar performance as other methods in high-entropy texts.
 2 Related Work
+In this section, we aim to review existing works on text watermarking algorithms.
+Overall, text watermarking can be divided into two different categories: watermarking for existing texts and watermarking for LLMs Liu et al.
+(2023).
+Text Watermarking for Existing Texts.
 Watermark message can be embedded into existing text by performing a series of modifications, such as lexical substitutions Munyer and Zhong (2023); Topkara et al.
@@
 (2023a) fine-tuned the infill model to generate consistent word distribution given several corrupted texts of the same origin.
-However, the level of modification to an existing text is often limited to protect text quality, thus restricting the capacity and effectiveness of lexical-based watermarks.
+However, the level of modification to an existing text is often limited to protect text quality in the watermarked output.
+Hence, watermarking for existing text is less favored due to its limited capacity and effectiveness.
+Text Watermarking for LLMs.
 Watermark embedding can also be conducted during the LLM generation phase, either by modifying the token sampling or model logits.
@@
 (2023b); Ren et al.
+(2023); He et al.
 (2024) and unforgeability Liu et al.
@@
 Liu et al.
-(2023b) proposed a different approach which utilizes the semantic meaning of previous generated text to determine the green lists.
+(2024b) proposed a different approach called SIR which utilizes the semantic meaning of previous generated text to determine the green lists.
+Inspired by Liu et al.
+(2024b), He et al.
+(2024) firstly proposed the Crosslingual Watermark Removal Attack (CWRA), then offered X-SIR as a defense method against CWRA.
 Also, to enhance the unforgeability of text watermark, Liu et al.
@@
 However, it requires a human code dataset to help manually determine the entropy threshold.
+To sum up, there lacks an effective method to improve the low-entropy detection problem.
 3 Preliminaries
-3.1 LLM
-Here we introduce the necessary concepts in this paper.
-A large language model (LLM) takes a prompt as input and outputs subsequent tokens as the corresponding response.
+3.1 Text Generation Process of LLMs
+Here we introduce the necessary concepts used in this paper.
+A LLM, $M$ takes a prompt as input and outputs subsequent tokens as the corresponding response.
 Specifically, we denote the initial prompt as $x^{prompt}$.
@@
 3.2 Text Watermarking
-Text watermarking is to embed a piece of message into the given text that can be decoded for the purpose of identification.
-Here we include a specific branch of it, defined as Watermark for LLMs by Liu et al.
-(2024), specifically the ones concerning Logits Generation.
-The general watermarking scheme used in this paper is called KGW Kirchenbauer et al.
-(2023a), which is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$.
-The modification on the original logits by adding a small value to the logits of green-list tokens will result in LLM’s preference on those tokens in the output text.
+The general watermarking scheme used in this paper is called KGW, which is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$.
+The modification on the original logits by adding a small value (the watermark strength $\delta$) to the logits of green-list tokens will result in LLM’s preference on those tokens in the output text.
 Thus such modification could likely produce a green token.
-The detection algorithm is to calculate the watermark level using the following equation
-$z=(|s|_{G}-\gamma T)/\sqrt{\gamma(1-\gamma)T},$ (1)
-where $|s|_{G}$ is the number of green tokens and $\gamma$ is the green-list ratio in the model vocabulary.
-The algorithm would output 1 if the value is significantly greater than a given threshold, and 0 otherwise.
-In this work, we only concern the watermark detection process.
+The detection algorithm is to calculate the watermark level using the following equation,
+$z=(|s|_{G}-\gamma|T|)/\sqrt{\gamma(1-\gamma)|T|},$ (1)
+where $|s|_{G}$ is the number of green tokens in text $T$ and $\gamma$ is the green-list ratio in the model vocabulary.
+The algorithm would output $Detect(T)=1$ if the value is greater than a given threshold, and $Detect(T)=0$ otherwise.
 3.3 Token Entropy and Low-entropy Scenario
-For the logits during each generation step, its degree of spread-out can be measured by calculating the entropy $SE(P_{LLM}(x^{prompt},T_{:(l-1)}))$.
-In this paper, we follow the spike entropy from Kirchenbauer et al.
-(2023a),
-$SE(k)=\sum_{v}{\frac{p_{v}}{1+zp_{v}}},$ (2)
-where $p$ is the logits when sampling the token $k$ and $z$ is a scalar.
-The entropy value achieves minimal if the logits concentrates on a single token and maximal given a uniform distribution.The watermark generation process is more effective on a high-entropy token where the logits distribution is nearly uniform, so that the modification could produce a green-list token easily.
-On the contrary, a low-entropy token is almost deterministic and hard to be modified.
+The entropy of a token is measured by the degree of spread-out of the logits generated by the LLM during the sampling process of that token. In this paper, we follow the spike entropy from KGW,
+$SE(k)=\sum_{v}{\frac{p_{v}}{1+\tau p_{v}}},$ (2)
+where $p_{v}$ is the logits value of each token $v$ in the vocabulary when sampling the token $k$, and $\tau$ is a scalar.
+The entropy value reaches the minimal value if the logits concentrates on a single token and the maximal value given a uniform distribution.
+The watermark generation process is more effective on a high-entropy token where the logits distribution is nearly uniform, so that the modification could produce a green-list token easily.
+On the other hand, a low-entropy token is almost deterministic and hard to be affected.
 Therefore, the green-list ratio for a low-entropy token is much lower, resulting in fewer green tokens in the output.
-A low-entropy scenario refers to the LLM-generated output consisting of a great number of low-entropy tokens.
-Here we denote the low-entropy token ratio to the whole text as $\varepsilon$.
-4 Method
-In this section, we will first explain the motivation by analysing the watermark effect on low-entropy tokens and traditional detection methods, then provide a detailed introduction to the proposed EWD.
+Therefore, low-entropy watermarked texts pose threats on the detection accuracy as they are easily identified as human-written.
+4 Proposed Method
+In this section, we will first explain the motivation by analysing the watermark generation effect on low-entropy tokens, then provide a detailed explanation of the proposed EWD.
 4.1 Motivation
-In KGW Kirchenbauer et al.
-(2023a)’s watermark generation method, a positive constant is added to the logits value of all green-list tokens so that the final output would prefer such tokens.
+Entropy plays an important role in the watermark generation process, yet its role in the detection process has been underestimated.
+In the KGW generator, a positive constant is added to the logits value of all green-list tokens.
 The distribution level of the logits, represented by the entropy, matters significantly during this process.
-When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this logits modification would very likely produce a green-list token.
-An average-high-entropy text would then consist of many green tokens and output a high z-score according to Eq.1.
-However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would hardly affect the token sampling and produce a green token with a much lower probability close to the green list ratio $\gamma$.
-As a result, the number of green tokens and the consequence detection z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1.
-This is because to the detector, tokens are of the same influence, no matter how difficult it is for the logits modification to affect them.
-4.2 Proposed Method
-In figure 1, we can observe that a high-entropy token is easier to be affected by the modified logits, and has a larger probability of being sampled from the green list.
-Since a low-entropy token is hard to affect, then whether it is sampled from the green list or not should have a limited influence on the overall watermark level.
-Specifically, we argued that when an easy-to-affect (high-entropy) token is not green, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-affect (low-entropy) token.
-That is to say, The influence of a red token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its entropy, and same for the influence of a green token to yielding a positive detection result:
-$I(Detect(T)=\textbf{1}(t\in G)|t)\propto SE(t).$ (3)
-When reflected in the real world, this influence index is represented by the token weight.
-Token with a larger spike entropy, meaning that it is easier to be affected, will be given a larger influence weight.
-On the contrary, a low-entropy token will be given a much lower weight close to 0.
-As the sum of all green token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a red low-entropy token than a red high-entropy token, and a larger gain when observing a green high-entropy token than a green low-entropy token.
+When the logits resembles a uniform distribution, this logits modification would very likely produce a green-list token.
+However, when the logits concentrates on one single token, this watermark embedding process would hardly affect the token sampling and produce a green token with a much lower probability close to the green list ratio $\gamma$.
+Since entropy determines the watermark generation process, then it should also affect the watermark detection.
+However, in the KGW detector, all tokens are of the same influence, no matter how their entropy varies.
+4.2 Entropy-based Text Watermarking Detection
+Our proposed EWD aims to fully consider the role of entropy in the watermark detection process.
+Specifically, we firstly propose a positive influence-entropy relation, then utilize entropy for detection weight customization.
+The influence of a token during detection should be proportional to its entropy.
+From previous sections, we explain why a high-entropy token is easier to be affected by the watermark generator, as well as the situation for a low-entropy token.
+Since a low-entropy token is hard to be affected, then whether it is sampled from the green list or not should have a limited influence on the overall watermark level.
+Similarly, when an easy-to-affect (high-entropy) token is not green, it should contribute more to showing that the text $T$ is not watermarked.
+Therefore, we introduce a positive influence-entropy relation to determine how much a token $t$ can influence the detection result:
+$I(Detect(T)=\textbf{1}(t\in G)|t)\propto SE(t),$ (3)
+where $T$ is the given text and $G$ is the green list.
+The detection weight of a token should be customized based on its entropy to fully reflect the positive influence-entropy relation.
+The influence index in Eq.
+(3) is represented by the detection weight of each token.
+However, the detection weight assigning of each token in SWEET fails to reflect the positive influence-entropy relation.
+This is because that inside the two groups of above-threshold tokens and below-threshold tokens, the weights of tokens with different entropy are identical.
+Therefore, to fully reflect the influence-entropy relation, we utilize a monotonically-increasing and continuous function $f$ to customize weights from token entropy.
+The weight $W(t)$ of a token $t$ can be illustrated as follows:
+$W(t)=f(SE(t)-C_{0}),$ (4)
+where $C_{0}$ is the minimal value of the spike entropy to normalize the entropy input before computing the weight.
+Using the entropy-based weight customization, even a sight increase in the token entropy will result in a increased detection weight.
+Table 1 summarizes the weight features of different methods and their reflection of the proposed influence-entropy relation.
+Table 1: Features of detection weights by different methods and how they can reflect the proposed positive influence-entropy relation.
+Methods Detection Weight
+Reflection of
+Eq. (3)
+KGW Constant None
+SWEET Binary Partial
+EWD Customized Full
 Algorithm 1 Entropy-based Watermark Detection
@@
      If current token $T_{l}$ is inside the green list $G$, then add its weight $W_{l}$ to $|s|_{G}$
-Compute the $z$-statistic:
-$\centering z=(|s|_{G}-\gamma\sum_{i=m}^{|T|-1}{W}_{i})/\sqrt{\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}}\@add@centering$ (4)
-if $z>\tau$ then return 1, $i.e.,$Input text is watermarked"
+Compute the detection score $z^{\prime}$ by Eq. (5)
+if $z^{\prime}>\tau$ then return 1, $i.e.,$Input text is watermarked"
 else return 0, $i.e.,$Input text is not watermarked"
-Now we are aware of the positive influence-entropy relation as suggest Eq.3.
-We decided to take a step further than SWEET: customizing influence weight according to token entropy to fully reflect the influence-entropy relation.
-While in SWEET, the weight of 1 for above-threshold tokens and the weight of 0 for below-threshold tokens partially reflect the relation, there is not reflection inside the entropy range of above-threshold tokens.
-Therefore, we utilized a monotonically-increasing and continuous function $f$ to generate weights from token entropy.
-The weight assigning function can be illustrated as follows:
-$W(t)=f(SE(t)-C_{0}).$ (5)
 Pseudocode of our proposed detection method is provided in Algorithm 1.
 We first acquire the text entropy by computing the model logits of each token.
-The next step is to determine the influence weight of each token, the output of a designed function with their entropy values as inputs.
-Due to the narrow range of entropy values, our goal here is to create a bigger gap between tokens of high entropy and low entropy.
-We achieve this by the function $ComputeWeight$, which is training-free and a totally automated process.
-Here, we first normalize the text entropy.
-Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights.
+The next step is to determine the influence weight of each token, the output of a designed function ComputeWeight with their entropy values as inputs.
 After obtaining the entropy-based weights for each token, we then apply the standard detection procedure by Kirchenbauer et al.
 (2023a) to determine the group of green tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens.
-Finally, we sum up the weights of the green tokens and calculate the z-statistic using Eq.4 given $|s|_{G}$ and weights of all detected tokens.
-The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
-In real-world low-entropy scenarios, With our EWD, the vast majority of low-entropy red tokens will not prevent the detector from yielding a positive conclusion.
-Meanwhile, facing human-written texts, any high-entropy red token will strongly suggest the detector to make a negative output.
+Finally, we sum up the weights of the green tokens $|s|_{G}$ and calculate the z-score by
+$\centering z^{\prime}=(|s|_{G}-\gamma\sum_{i=m}^{|T|-1}{W}_{i})/\sqrt{\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}},\@add@centering$ (5)
+given $|s|_{G}$ and weight $W_{t}$ of each detected token $t$ in text $T$.
+The detector would give a positive result if the z-score is greater than the given threshold.
+It is noticably that the proposed algorithm is a fully automated process that eliminates the need for manual operations from the SWEET method.
 5 Theoretical Analysis
-In this section, we theoretically analyzed the Type-I and Type-II error of our proposed detection method. Specifically, we compared EWD with two previous detection methods, KGW and SWEET.
-To do so, we adopted Kirchenbauer et al.
-(2023a)’s steps of approach.
-Each token in both watermarked and non-watermarked texts can either be green or not, so we can model it as a binomial distribution.
+In this section, we theoretically analyze the Type-I and Type-II error of our proposed detection method. Specifically, we compare EWD with two previous detection methods, KGW and SWEET.
+To do so, we adopt the same steps as in KGW Kirchenbauer et al.
+(2023a).
+Each token in both watermarked and non-watermarked texts can either be green or red, so we can model it as a binomial distribution.
 While Zhao et al.
-(2023) set window size in the watermark generator as zero (one universal green list for all tokens) and exhibited major token dependence, $i.e.$, the same tokens in the text are of the same color (either all sampled from the green list or all from the red list), in our case, where a non-zero window size is adopted, the token dependence is significantly weakened.
+(2023) set the window size in the watermark generator as zero (one universal green list for all tokens) and exhibits major token dependence, $i.e.$, the same tokens in the text are of the same color, in our case, where a non-zero window size is adopted, the token dependence is significantly weakened.
 For example, the same token with different prefix would exhibit different sampling result.
-Therefore, to simplify and facilitate the process of theoretical analysis, we made the reasonable assumption that the sampling of each token is independent.
+Therefore, to simplify and facilitate the process of theoretical analysis, we make the reasonable assumption that the sampling of each token is independent.
 Hence, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes.
@@
 Type-I error measures the likelihood a human-written text is incorrectly identified as being generated by a watermark algorithm.
-In a human-written text, each token is independent of the watermark algorithm.
+In a human-written text $T$, each token is independent of the watermark algorithm.
 Therefore, its probability to be included in the green list is $\gamma$.
-Therefore, KGW Kirchenbauer et al.
-(2023a)’s sum of green tokens is approximated by $\mathcal{N}(\gamma T,\gamma(1-\gamma)T)$.
-For SWEETLee et al.
-(2023) that only considers high-entropy tokens, such distribution is modelled as $\mathcal{N}(\gamma\tilde{T},\gamma(1-\gamma)\tilde{T})$, where $\tilde{T}$ is the text length excluding low-entropy tokens.
-Our method EWD assigned weights to tokens based on entropy, leading to a weighted binomial distribution.
+Therefore, KGW’s sum of green tokens is approximated by $\mathcal{N}(\gamma|T|,\gamma(1-\gamma)|T|)$.
+For SWEET that only considers high-entropy tokens, such distribution is modelled as $\mathcal{N}(\gamma\tilde{|T|},\gamma(1-\gamma)\tilde{|T|})$, where $\tilde{T}$ is the text excluding low-entropy tokens.
+Our method EWD assign weights to tokens based on entropy, leading to a weighted binomial distribution.
 We could also approximate it to a normal distribution with mean $\mu=\gamma\sum_{i=m}^{|T|-1}{W}_{i}$ and variance $\sigma^{2}=\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}$.
-When $z$ is set to 2, the probabilities for $|s|_{G}$ under each method to surpass its detection threshold $|\tilde{s}|_{G}$ is identically 2.28%.
+The theoretical Type-I errors for all detection methods are the same.
+When the detection threshold is set to 2, the probabilities for $|s|_{G}$ under each method to surpass its detection threshold $|\tilde{s}|_{G}$ is identically 2.28%.
+5.2 Type-II Error
+Type-II error measures the likelihood a machine-generated text is incorrectly identified as human-written.
+For analysis, we assume that a sequence generated by the watermark algorithm has a length of $|T|=200$ tokens and the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
+Originally, KGW samples over 500 watermark-generated news scripts and reports an average spike entropy of 0.807.
+In this section, we focus specifically on cases with a much lower average spike entropy.
+Based on our data, we approximate the distribution of spike entropy in low-entropy watermarked texts with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$.
+Under this distribution, the average spike entropy is 0.608.
+The following probability from KGW shows a lower bound of the probability of a token $k$ being sampled from the green list:
+$\mathbb{P}[k\in G]\geq\frac{\gamma\alpha}{1+(\alpha-1)\gamma}SE(k),$ (6)
+where $\alpha=exp(\delta)$. We would utilize this probability later in the calculation of $\mu$ and $\sigma^{2}$.
+KGW.
+Based on Eq.
+(6), the statistical $\mu$ and upper bound of $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]$ and $\sum_{k}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$.
+When the average entropy $SE^{*}$ given by the simulated power law distribution is 0.608, we approximate $|s|_{G}$ by $\mathcal{N}(107.10,49.70)$.
+If we set a threshold of $z$ = 2 (corresponding to $|\tilde{s}|_{G}$ = 114.14) for detection, the false negative rate exceeds 84.1%.
+This means that KGW is not good at detecting the watermarked texts in cases of low entropy.
+SWEET.
+Following SWEET, we set the entropy threshold at 0.695 and exclude all tokens below this threshold during detection.
+The average spike entropy $SE^{*}$ and remaining text length $\tilde{|T|}$ then change to 0.82 and 24, respectively.
+Using the equations from KGW, we can get the approximation as $\mathcal{N}(17.48,4.84)$.
+If we set a threshold of $z$ = 2 (corresponding to $|\tilde{s}|_{G}$ = 17.02) for detection, the false negative rate exceeds 41.7%.
+EWD.
+EWD assigns higher influence weights to tokens with high spike entropy.
+Note that in experiments, we use several weight functions, and here, for analysis purposes, we use a linear weight function for each token $k$: $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant.
+Based on Eq.
+(6), the statistical $\mu$ and $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ and $\sum_{k}{W(k)}^{2}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$.
+Therefore, we can approximate the $|s|_{G}$ detected by EWD with $\mathcal{N}(5.79,0.31)$.
+If we set a threshold of $z^{\prime}$ = 2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.4% (details in Appendix A).
+Table 2 summarizes our calculated theoretical Type-I and Type-II error using different methods in low-entropy detection scenarios.
+Overall, our method achieves the lowest Type-II error among all methods while maintaining the same Type-I error.
+Table 2: This table summarizes our calculated theoretical Type-I and Type-II error using different methods in low-entropy detection scenarios.
+The data in the table is presented in percentages (%).
+Lower values indicates better detection accuracy.
+Methods Type-I Error Type-II Error
+KGW 2.28 84.1
+SWEET 2.28 41.7
+EWD 2.28 33.4
+6 Experiments
 Table 3: Main results on different datasets using KGW, SWEET and our EWD for detection.
@@
 EWD 1.000 0.997 1.000 0.978 1.000 0.995 0.993 1.000 0.976 0.997
-5.2 Type-II Error
-Type-II error measures the likelihood a machine-generated text is incorrectly identified as human-written.
-For analysis, we assumed that a sequence generated by the watermark algorithm has a length of $T=200$ tokens and the watermark algorithm has hyper-parameters $\gamma=0.5$ and $\delta=2$.
-Originally, Kirchenbauer et al.
-(2023a) sampled over 500 watermark-generated news scripts and reported an average spike entropy of 0.807.
-In this section, we focused specifically on cases with a much lower average spike entropy.
-Based on our data, we approximated the distribution of spike entropy in low-entropy watermarked texts with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$.
-Under this distribution, the average spike entropy is 0.607.
-The following probability from Kirchenbauer et al.
-(2023a) showed a lower bound of the probability of a token $k$ being sampled from the green list:
-$\mathbb{P}[k\in G]\geq\frac{\gamma\alpha}{1+(\alpha-1)\gamma}SE(k),$ (6)
-where $\alpha=exp(\delta)$. We would utilize this probability later in the calculation of $\mu$ and $\sigma^{2}$.
-KGW.
-Based on Eq.6, the statistical $\mu$ and upper bound of $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]$ and $\sum_{k}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$.
-When the average entropy $SE^{*}$ given by the simulated power law distribution is 0.607, we approximated $|s|_{G}$ by $\mathcal{N}(106.9,7.1)$.
-If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 114) for detection, the false negative rate exceeds 84.6%.
-This means that KGW is not good at detecting the watermarked texts in cases of low entropy.
-SWEET.
-Following SWEET, we set the entropy threshold at 0.695 and excluded all tokens below this threshold during detection.
-The average spike entropy $SE^{*}$ and remaining text length $\tilde{T}$ then changed to 0.82 and 24 respectively.
-Using the equations from KGW, we can get the approximation as $\mathcal{N}(17.5,2.2)$.
-If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 17.2) for detection, the false negative rate exceeds 42%.
-EWD.
-We assigned higher influence weights to tokens with high spike entropy in order to achieve a better detection rate in cases of low average spike entropy.
-Note that in experiments, we use several weight functions, and here, for analysis purposes, we used a linear weight function for each token $k$: $W(k)=SE(k)-C_{0}$, where $C_{0}$ is a constant.
-Based on Eq.6, the statistical $\mu$ and $\sigma^{2}$ can be obtained as $\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ and $\sum_{k}{W(k)}^{2}\mathbb{P}[k\in G]\cdot(1-\mathbb{P}[k\in G])$.
-Therefore, we can approximate the $|s|_{G}$ detected by EWD with $\mathcal{N}(5.79,0.56)$.
-If we set a threshold of $z$=2 (corresponding to $|\tilde{s}|_{G}$ = 5.55) for detection, the false negative rate exceeds 33.3% (Proof and details in Appendix A).
-Overall, Our method has a type-II error much lower than KGW and SWEET.
-6 Experiments
+In this section, we first introduce the tasks and datasets used for watermarked text generation.
+Then, we compare the performance of our EWD against several baselines on those datasets.
+Next, we analyze the role of our proposed entropy-based weight and study the performance of EWD using different weight computing functions.
+Finally, we evaluate the detection performance without the original prompt, detection speed and robustness of our EWD.
 6.1 Experiment Settings
-Figure 2: The left figure shows the z-scores of watermarked and human texts in the Rotowire, HumanEval and MBPP datasets respectively, each being detected with 3 different methods.
-The right figure shows the relationship between token weights and the probability of being sampled from the green list in both watermarked and human texts.
-Figure 3: We utilized a variety of weight functions and measured their performance in the code detection datasets.
-The left figure is a visualization of all studied functions, with normalized spike entropy as input.
-The right figure shows each function’s detection F1 under 1% FPR with comparison to the SWEET baseline.
+Figure 2: Subfigure (a) shows the z-scores of watermarked and human texts in the Rotowire, HumanEval and MBPP datasets, respectively, each being detected with 3 different methods.
+Subfigure (b) shows the relationship between token weights and the probability of being green in both watermarked and human texts.
+Figure 3: We utilize two additional weight functions other than Linear and measured their performance in the code detection datasets.
+Subfigure (a) is a visualization of all studied functions, with normalized spike entropy as input.
+Subfigure (b) shows each function’s detection F1 under 1% FPR with comparison to the SWEET baseline.
 Each data point can correspond to the illustrated function on the left figure by looking at its shape and x-axis value.
-Tasks and Datasets. First of all, we used the KGW Kirchenbauer et al. (2023a) to embed watermark in all the following tasks.
-To evaluate the detection performance of our proposed method, we targeted at a commonly-seen low-entropy scenario, code generation.
-We adopted two datasets following Lee et al.
+Tasks and Datasets.
+There are various text generation tasks in the field of Natural Language Processing, and here we aim to generate watermarked text of both low-entropy and high-entropy distributions.
+Since our focus is on comparing different detection methods, we utilize the same KGW watermark generation algorithms in all the following tasks and datasets.
+For low-entropy scenarios, we target at the task of code generation by adopting two datasets following Lee et al.
 (2023), HumanEval Chen et al.
@@
 (2021).
-We used StarCoder Li et al.
-(2023) for code generation.
-And we used the human-written answers as non-watermarked texts.
-During evaluation, we restricted the minimal length of both samples to be 15 tokens.
-Besides, we fine-tuned a T5-small Raffel et al.
-(2023) model for data-to-text generation.
-Specifically, we utilized Wu et al.
+The two datasets contain test cases of python programming problems and corresponding reference answers, which are utilized as human-written samples.
+Code generation is conducted by StarCoder Li et al.
+(2023).
+During evaluation, the length of both samples are restricted to be at least 15 tokens.
+The first common high-entropy scenario is the news report generation.
+Our input prompts to the pre-trained OPT-1.3B Zhang et al.
+(2022) are from the C4 Raffel et al.
+(2020) dataset.
+Besides, we include another high-entropy data-to-text generation task by utilizing Wu et al.
 (2021)’s version of the Rotowire dataset Wiseman et al.
-(2017), which was converted for a text-to-table task.
+(2017).
 The dataset consists of NBA game records, presented as tables, and human-written descriptions.
-We used the table information as input and generate descriptive texts.
-For each input prompt, the models are expected to generate 200 ± 5 tokens.
-We also utilized the pre-trained OPT-1.3B Zhang et al.
-(2022) for news report generation.
-Our input prompts are from C4 Raffel et al.
-(2020) dataset.
-We restricted the maximum prompt length to 50 tokens.
-For each input prompt, the models are expected to generate 200 ± 5 tokens using binomial sampling and beam-search.
+We fine-tune a T5-small Raffel et al.
+(2023) model for text generation and input the table information to the fine-tuned model.
+We generate 200 samples of length 200 ± 5 tokens for each high-entropy tasks, using binomial sampling and beam search.
 Baselines and Evaluation Metrics.
-We selected two detection methods as baselines.
-The first is KGW detector Kirchenbauer et al.
-(2023a).
-The second is SWEET’s detector Lee et al.
-(2023), which excludes low-entropy tokens during detection.
-We follow the instructions in SWEET to select entropy threshold from the median, mean, and 3rd quartile (Q3) values of the entropy computed from human texts.
+Two detection methods are selected as baselines.
+The first is KGW detector.
+The second is SWEET’s detector, which excludes low-entropy tokens during detection.
+Note that SWEET proposes both selective watermark generation and detection, while here we only adopt the detection part.
+The selection of the entropy threshold follows the instructions in SWEET.
 For evaluation, we set the false positive rate (FPR) under 5%, following Zhao et al.
 (2023)’s instruction.
-We reported the true positive rate (TPR) and F1 score for each method.
+We report the true positive rate (TPR) and F1 score for each method.
 Hyper-parameters.
-For watermarked algorithm, we set $\gamma$ and $\delta$ to 0.5 and 2 respectively.
-Binomial sampling set the sampling temperature to 0.7.
-Beam-search set the number of beams to 8.
-To avoid excessive repetition of text, we set "no_repeated_ngrams" to 16.
+For the KGW generator, $\gamma$ and $\delta$ is set to 0.5 and 2, respectively.
+Binomial sampling sets the sampling temperature to a common value of 0.7.
+Beam-search sampling sets the number of beams to 8.
+To avoid excessive repetition of text, "no_repeated_ngrams" is set to 16.
 6.2 Main Results
-Before detection, we measured the generation quality of all generated texts.
+Before detection, we measure the generation quality of all generated texts.
 Specifically, we are able to achieve 31.0% and 30.4% in pass@1Chen et al.
-(2021) for HumanEval and MBPP dataset respectively, comparable to Lee et al.
+(2021) for the HumanEval and MBPP dataset, respectively, comparable to Lee et al.
 (2023)’s.
@@
 Table 3 shows that for the low-entropy scenario, watermarked code detection, our detector would achieve the best detection accuracy against other baselines.
-Specifically, for HumanEval and MBPP respectively, our method could achieve a 3.5% and 9.4% improvement than SWEET in true positive rate while maintaining a low-than 5% false positive rate.
-Table 1 shows that for the high-entropy scenarios using Rotowire and C4 datasets, our method is overall very similar to other baselines, sometimes even better than them.
-6.3 Analysis
+Specifically, for HumanEval and MBPP, respectively, our method could achieve a 2.5% and 9.4% improvement than SWEET in TPR while maintaining a low-than 5% FPR.
+For the high-entropy scenarios using the Rotowire and C4 datasets, the detection performance of our method is overall very similar to other baselines, sometimes even better than some baselines.
+6.3 Empirical Analysis
 Impact of Entropy-based Weights.
-In figure 2, we gave a few visualizations to show the reason why EWD surpassed other baselines in detection performance.
+In Figure 2, we provide a few visualizations to show the reason why EWD surpasses other baselines in detection performance.
 Figure 2(a) shows the z-scores of both watermarked and human texts in different datasets, detected by different methods.
-Our EWD method would result in overall higher z-score for watermarked texts, and slightly lower z-score for human texts.
+Our EWD method would result in overall higher z-scores for watermarked texts, and slightly lower z-scores for human texts.
 The z-score increase is more significant in low-entropy code generation tasks than in the Rotowire dataset.
 By enlarging the gap between the z-scores of watermarked and human texts, our detector could better distinguish the watermarked texts.
-In figure 2(b), we plotted the relationship of a token’s entropy-based weight and probability of that token being sampled from the green list.
-In watermarked text, the probability is proportional to the token weight.
-That is to say, the tokens assigned with a large weight are very likely to be sampled from the green list and contribute to a larger z-score.
-On the contrary, in human-written texts, there doesn’t exist such a relationship, explaining why the z-scores of human text remain low.
-Ablation Study.
-In the experiments, we utilized a linear function to gain the entropy-based weights for each token.
-We also studied a variety of monotonic increasing functions as shown in figure 3, which are applied in Eq.5 as $f$.
+In Figure 2(b), we plot the relationship of a token’s entropy-based weight and probability of that token being sampled from the green list.
+In human-written texts, there does not exist a relationship between the probability and token weights, explaining why the z-scores of human text remain low.
+Meanwhile, in watermarked texts, the probability is proportional to the token weight, which means the tokens assigned with a large weight are very likely to be sampled from the green list and contribute to a larger z-score.
+This proves that our entropy-based weights can help more accurately reflect the watermark level in low-entropy texts.
+Performance with Different Weight Functions.
+In the experiments, we utilize a linear function to gain the entropy-based weights for each token.
+We also study two more monotonic increasing functions, the Sigmoid and the Exponential function.
+As shown in Figure 3(a), we further change the function strength to 4 different values in each function, which represents how far the function deviates from being linear, for example, "Sigmoid 10" deviates from linear more than "Sigmoid 8".
+The abovementioned functions are applied in Eq.
+(4) as $f$.
 Specifically, each function’s gradient is different.
-For our custom Sigmoid functions, the gradient starts off high and gradually decreases, that is to say, tokens with relatively low entropy will have significant differences in weights, while tokens with relatively high entropy will exhibit smaller differences.
-Similarly for our custom Exponential functions, the gradient starts off low and gradually increases, meaning that tokens with relatively low entropy will have smaller differences in weight, while tokens with relatively high entropy will exhibit larger differences.
-We use function strength to represent how far the function deviates from being linear, for example, "Sigmoid 10" deviates from linear more than "Sigmoid 8".
-In the experiments with those functions and two code generation datasets, we can observe that the total 8 listed functions could surpass the SWEET baseline in terms of detection F1.
-This reflects the versatility of our method, as it is applicable to a wide range of functions.
-Time Complexity.
-Table 2: The table below shows the average time taken to generate 200-token texts using KGW generator, as well as the average time taken for detection using different methods, measured in seconds (s).
+For our custom Sigmoid functions, the gradient starts off high and gradually decreases, meaning that tokens with relatively low entropy will have significant differences in weights, while tokens with relatively high entropy will exhibit smaller differences.
+Similarly, for our custom Exponential functions, the gradient starts off low and gradually increases, meaning that tokens with relatively low entropy will have smaller differences in weight, while tokens with relatively high entropy will exhibit larger differences.
+Figure 3(b) demonstrates that the two listed functions with different strength could surpass the SWEET baseline in terms of detection F1.
+This reflects the versatility of our method, as it is applicable to more functions.
+Table 4: Detection performance of watermarked code using the general prompt instead of original prompts.
+Detection w/ General Prompt
+Methods F1 under 5%FPR Best F1
+KGW 0.498 0.741
+SWEET 0.580 0.762
+EWD 0.600 0.764
+[…]
+Detection Speed.
+Table 5: This table shows the average time taken to generate 200-token texts using KGW generator, as well as the average time taken for detection using different methods, measured in seconds.
 The experiment is conducted on a NVIDIA RTX A6000 GPU.
@@
 Compared to the KGW detection, both SWEET and our proposed EWD would demand extra time to compute token entropy.
-Specifically, our detector would also compute weight given the token entropy before calculating the final z-score.
-According to table 2, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our EWD respectively.
+Specifically, our detector would also compute weights given the token entropy before calculating the final z-score.
+According to Table 5, the detection time per text is 0.0391, 0.0830 and 0.0831 seconds for KGW, SWEET and our EWD, respectively.
 In spite of being double the value of KGW, our detector remains highly efficient, and the difference with SWEET is virtually insignificant.
-Performance against Attacks.
+Performance against the Back-translation Attack.
 Table 6: Detection performance of back-translated watermarked texts using different detection methods.
@@
 EWD 0.900 0.943 0.942 0.947
-Watermarked texts usually would be edited, which would cause detection performance drop.
-Here we also evaluated the baseline methods and our EWD under this setting by utilizing back-translation as the attack to remove watermark.
-Specifically, we firstly generated watermarked texts using the C4 dataset, then translated them from English to French, and later back to English again for detection.
-In table 3, we can observe that our EWD would achieve the best detection accuracy under 1% FPR and similar results with KGW under 5% FPR.
-[…]
-Table 4: Detection performance of watermarked code using the general prompt instead of original prompts.
-Detection w/ General Prompt
-Method F1 under 5%FPR Best F1
-KGW 0.498 0.741
-SWEET 0.580 0.762
-EWD 0.600 0.764
+Watermarked texts usually would be edited before being detected, which would remove part of the watermark and cause detection performance drop.
+Here we also evaluate the baseline methods and our EWD under this setting by utilizing back-translation as the attack to remove watermark.
+Specifically, we firstly generate watermarked texts using the C4 dataset, then translate them from English to French, and later back to English again for detection.
+In Table 6, our EWD would achieve the best detection accuracy under 1% FPR and similar results with KGW under 5% FPR.
 7 Conclusion
-In this work, we proposed a text watermark detection algorithm called EWD based on token entropy.
-We gave theoretical analysis on the detection properties of our method in low-entropy scenarios with comparison of previous works.
-We also validated the detection performance of our method in the low-entropy code generation task with a number of weight functions.
-Furthermore, our method would achieve similar performance in high-entropy scenarios as the previous works.
-Experiments also showed that compared to previous methods, our method would not bring extra computational cost.
+In this work, we propose a text watermark detection method called EWD that fully considers the influence of token entropy.
+We first establish the positive influence-entropy relation in the detection process, then utilize entropy for weight customization to fully reflect the proposed relation.
+We provide theoretical analysis on the detection properties of different methods in low-entropy scenarios.
+Experiments validate the detection performance of our method in the low-entropy code generation task with a number of weight functions.
+Furthermore, in high-entropy scenarios, our method would achieve similar performance as previous works in detection accuracy, detection speed and robustness.
 Limitations
-The limitations of our method mainly include two parts.
-The first part is that the low-entropy datasets tested are limited.
-We only conducted experiments on two code generation datasets, and we did not conduct experiments on more low-entropy datasets.
-At the same time, our method should be theoretically effective for various types of watermark methods, such as token sampling-based methods.
-This paper did not try more types of methods, which is an important future work direction.
+Our method mainly includes two limitations.
+The first one is that the low-entropy datasets tested are limited.
+We utilize two code generation datasets for evaluation.
+In the future, we aim to include more low-entropy tasks and datasets.
+Secondly, our method should be theoretically effective for various types of watermark methods, such as token sampling-based methods.
+In the future, we aim to implement our method on more types of text watermarking frameworks.
+Acknowledgement
+We sincerely appreciate the insightful comments from all reviewers. Their help has greatly contributed to the improvement of this work.
+The work described in this paper was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 14222922, RGC GRF 2151185).
 Appendix A Proof of Type-II Error for EWD
-Here we demonstrate how we obtained the theoretical Type-II error for EWD.
-First of all, we attempted to calculate the theoretical $\mu$ and $\sigma^{2}$ based on Eq.6.
+Here we demonstrate how we obtain the theoretical Type-II error for EWD.
+First of all, we attempt to calculate the theoretical $\mu$ and $\sigma^{2}$ based on Eq.
+(6).
 The theoretical $\mu$ is the weighted sum of green tokens in the watermarked text, using ${|s|_{G}}_{k}$ to represent the weighted number of green tokens from each token $k$,thus we can obtain:
-$\mu$ $=\sum_{k}{|s|_{G}}_{k}$ (7)
-$\mu$ $=\sum_{k}\mathbb{P}[k\in G]\cdot W(k)$ (8)
-$=T\mathbb{E}\frac{\gamma\alpha SE}{1+(\alpha-1)\gamma}(SE-C_{0})$ (9)
+$\mu$ $=\sum_{k}{|s|_{G}}_{k},$ (7)
+$=\sum_{k}\mathbb{P}[k\in G]\cdot W(k),$ (8)
+$=T\mathbb{E}\frac{\gamma\alpha SE}{1+(\alpha-1)\gamma}(SE-C_{0}),$ (9)
 $=T\frac{\gamma\alpha}{1+(\alpha-1)\gamma}(\mathbb{E}[SE^{2}]-C_{0}\mathbb{E}[SE]).$ (10)
-Given the simulated power law of spike entropy, we can obtain $E[SE^{2}]$ and $E[SE]$ as 0.377 and 0.608 respectively, producing $\mu$ equal to 5.79.
+Given the simulated power law of spike entropy, we can obtain $E[SE^{2}]$ and $E[SE]$ as 0.377 and 0.608, respectively, producing $\mu$ equal to 5.79.
 And for $\sigma^{2}$, we first obtain the variance for each token $k$:
-${\sigma_{k}}^{2}$ $=\mathbb{E}[({|s|_{G}}_{k}-\mathbb{E}[{|s|_{G}}_{k}])^{2}]$ (11)
-$=\mathbb{E}[{|s|_{G}}_{k}^{2}]-\mathbb{E}[{|s|_{G}}_{k}]^{2}$ (12)
-$=\mathbb{P}[k\in G]\cdot W(k)^{2}-\mathbb{P}[k\in G]^{2}\cdot W(k)^{2}$ (13)
+${\sigma_{k}}^{2}$ $=\mathbb{E}[({|s|_{G}}_{k}-\mathbb{E}[{|s|_{G}}_{k}])^{2}],$ (11)
+$=\mathbb{E}[{|s|_{G}}_{k}^{2}]-\mathbb{E}[{|s|_{G}}_{k}]^{2},$ (12)
+$=\mathbb{P}[k\in G]\cdot W(k)^{2}-\mathbb{P}[k\in G]^{2}\cdot W(k)^{2},$ (13)
 $=W(k)^{2}\cdot(\mathbb{P}[k\in G])(1-\mathbb{P}[k\in G]).$ (14)
 Therefore, if we represent $\frac{\gamma\alpha}{1+(\alpha-1)\gamma}$ by $C_{1}$, the variance of the whole text can be formulated as:
-$\sigma^{2}$ $=\sum_{k}W(k)^{2}\cdot(\mathbb{P}[k\in G])(1-\mathbb{P}[k\in G])$ (15)
-$=T\mathbb{E}[W(k)^{2}C_{1}SE(1-C_{1}SE)]$ (16)
-$=TC_{1}\mathbb{E}[(SE-C_{0})^{2}SE(1-C_{1}SE)]$ (17)
+$\sigma^{2}$ $=\sum_{k}W(k)^{2}\cdot(\mathbb{P}[k\in G])(1-\mathbb{P}[k\in G]),$ (15)
+$=T\mathbb{E}[W(k)^{2}C_{1}SE(1-C_{1}SE)],$ (16)
+$=TC_{1}\mathbb{E}[(SE-C_{0})^{2}SE(1-C_{1}SE)],$ (17)
 $=TC_{1}(-C_{1}\mathbb{E}[SE^{4}]+(2C_{0}C_{1}+1)\mathbb{E}[SE^{3}]$
 $\ -(2C_{0}+{C_{0}}^{2}C_{1})\mathbb{E}[SE^{2}]+{C_{0}}^{2}\mathbb{E}[SE]).$ (18)
-Given the simulated power law of spike entropy, we can obtain $E[SE^{4}]$ and $E[SE^{3}]$ as 0.24 and 0.158 respectively, producing $\sigma$ equal to 0.56.
-When the detection threshold $z$ is set as 2 (corresponding to $|\tilde{s}|_{G}$ = 5.55), the probability of $|s|_{G}\sim\mathcal{N}(5.79,0.56)$ falling below 5.55 is 33.3%, $i.e.$ Type-II error.
+Given the simulated power law of spike entropy, we can obtain $E[SE^{4}]$ and $E[SE^{3}]$ as 0.24 and 0.158, respectively, producing $\sigma$ equal to 0.56.
+When the detection threshold $z^{\prime}$ is set as 2 (corresponding to $|\tilde{s}|_{G}$ = 5.55), the probability of $|s|_{G}\sim\mathcal{N}(5.79,0.31)$ falling below 5.55 is 33.4%, $i.e.$, the Type-II error.